### v1

# Query-Specific GNN: A Comprehensive Graph Representation Learning Method for Retrieval Augmented Generation

Abstract
Retrieval-augmented generation (RAG) has demonstrated its ability to enhance Large Language Models (LLMs) by integrating external knowledge sources. However, multi-hop questions, which require the identification of multiple knowledge targets to form a synthesized answer, raise new challenges for RAG systems. Under the multi-hop settings, existing methods often struggle to fully understand the questions with complex semantic structures and are susceptible to irrelevant noise during the retrieval of multiple information targets. To address these limitations, we propose a novel graph representation learning framework for multi-hop question retrieval. We first introduce a Multi-information Level Knowledge Graph (Multi-L KG) to model various information levels for a more comprehensive understanding of multi-hop questions. Based on this, we design a Query-Specific Graph Neural Network (QSGNN) for representation learning on the Multi-L KG. QSGNN employs intra/inter-level message passing mechanisms, and in each message passing the information aggregation is guided by the query, which not only facilitates multi-granular information aggregation but also significantly reduces the impact of noise. To enhance its ability to learn robust representations, we further propose two synthesized data generation strategies for pre-training the QSGNN. Extensive experimental results demonstrate the effectiveness of our framework in multi-hop scenarios, especially in high-hop questions the improvement can reach 33.8%. The code is available at: https://github.com/Jerry2398/QSGNN.
1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a powerful paradigm for enhancing the capabilities of Large Language Models (LLMs)  (Peng et al., 2024; Procko & Ochoa, 2024; Gao et al., 2023). By retrieving pertinent information from external knowledge sources and integrating it into the generation process, RAG enables LLMs to ground their responses in factual information, significantly reducing the propensity for hallucinations  (Fan et al., 2024).
A challenging scenario in RAG is handling multi-hop questions, where the answer cannot be directly supported by a single document [^1] Following prior works  (Edge et al., 2024; Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025), we view document as the retrieval target in this paper. but requires information synthesized from multiple interrelated documents, each contributing a unique piece of the puzzle for the answer  (Dua et al., 2019; Chen et al., 2019). These questions demand that the RAG system accurately identify multiple relevant documents and synthesize them to form a coherent answer. Traditional RAG approaches, which primarily rely on semantic similarity for retrieval  (Karpukhin et al., 2020; Chen et al., 2024; Jiang et al., 2023; Trivedi et al., 2022b), often struggle with multi-hop questions since they overlook the relational dependencies between documents. To overcome this limitation, recent efforts have proposed to leverage Knowledge Graphs (KGs). These methods first construct KGs to model the relationships between documents. During retrieval stage, they either navigate the KGs using graph search strategies  (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2025) or utilize Graph Neural Networks (GNNs)  (Fang et al., 2019; Mavromatis & Karypis, 2024; Luo et al., 2025) to identify relevant information. By using KGs, these approaches can better capture the dependencies between documents, thereby improving the retrieval process for multi-hop questions.
Figure 1: Performance on multi-hop questions. Existing methods struggle to perform well as hop number increases while QSGNN can achieve better performance on high-hop questions.
Although KG based methods have shown improved performance over traditional approaches, they still face significant challenges in addressing the complexities of multi-hop questions. These challenges can be summarized as follows: i) Semantic Comprehensiveness. Multi-hop questions are inherently more complex than one-hop questions from the semantic perspective, requiring a comprehensive understanding of diverse information components. For instance, consider the question: “What was Professor Joe’s opinion about the Apple event in 2023 that featured the announcement of a foldable smartphone concept?” In this case, the entity “Apple” has multiple possible meanings (e.g., a company or a fruit), and the phrase “announcement of a foldable smartphone concept” may refer to a sentence or even a paragraph whose meaning cannot be fully captured by discrete entities within the KG. Such questions not only require the retrieval process to understand context-dependent entities but also demand a full understanding of multi-granular information. However, existing methods primarily focus on entities while ignoring the contextual and multi-granular information. This limitation may lead to the inclusion of irrelevant information (e.g., descriptions of apples as fruits) or the loss of critical details (e.g., the absence of precise entities like “foldable smartphone concept” in the KG), ultimately compromising the performance. ii) Noise Sensitivity. The retrieval process for multi-hop questions is highly sensitive to noise since it has multiple retrieval targets. The noise in any one retrieval target will disrupt the final result. However, existing methods are mostly prone to noise influence. For instance, graph search based methods rely on heuristic strategies to explore relevant information, which is likely to incorporate irrelevant nodes or miss important information. GNN based methods use message passing, which may also aggregate information from irrelevant neighbors. Furthermore, the large search space of KG exacerbates this issue. As shown in Figure  1, the performance of existing methods degrades significantly for high-hop questions, because the number of irrelevant nodes in the KG increases exponentially as the retrieval step increases.
To address these limitations, we propose an effective graph representation learning framework to enhance the RAG performance on multi-hop questions. Our approach is built on three key innovations: i) To comprehensively capture multiple information granularities and their complex relationships, we propose a Multi-information Level Knowledge Graph (Multi-L KG). It consists of three node levels, each representing distinct information granularity. The entity-level captures the basic semantic relationships, the chunk-level models the local contextual relationships, and the document-level represents the global thematic relationships. This design encompasses the fine-grained information relationships and facilitates a more comprehensive understanding for multi-hop questions. ii) To learn representations from the Multi-L KG for multi-hop retrieval, we design a novel GNN model called Query-Specific Graph Neural Network (QSGNN). Specifically, it adopts two kinds of message passing to handle the multiple information components of Multi-L KG: intra-level and inter-level message passing. Intra-level message passing focuses on the basic semantic relationships and the logical coherence relationships within each level, while inter-level message passing considers local-global relationships across levels. In each type of message passing, it considers both semantic relationships and alignment with the query. This design integrates multi-level information and significantly reduces the influence of irrelevant nodes, resulting in high-quality representations for multi-hop retrieval. iii) To further enhance the capability of QSGNN, we propose two data generation strategies to produce synthesized QA pairs for pre-training. This process requires no extra human cost and can be conducted after the construction of Multi-L KG. After pre-training, QSGNN can be quickly adapted to downstream tasks. We conduct extensive experiments to evaluate our method, which demonstrate that QSGNN achieves state-of-the-art performance compared to existing methods, especially in high hop questions the improvement can reach 33.8%.
2 Related Work
Retrieval-augmented generation(RAG). RAG aims at enhancing large language models (LLMs) by retrieving relevant information from external knowledge sources and integrating it into the generation process to enable fact-grounded responses  (Gao et al., 2023; Fan et al., 2024). Traditional RAG methods  (Chen et al., 2024; Li et al., 2023) encode documents from external knowledge sources into dense embeddings and retrieve relevant information by calculating similarity between query and document embeddings. However, these approaches fail to capture the complex relationships between documents, struggling to perform well on the multi-hop questions. In order to adapt the RAG to multi-hop questions, subsequent works  (Jiang et al., 2023; Trivedi et al., 2022b; Su et al., 2024) have proposed multi-step retrieval strategies, where they decompose the original question into multiple sub-questions via LLMs and retrieve relevant information for each sub-question. Although these methods explicitly model the internal structure of the query, they still neglect the multi-level relationships between documents, resulting in limited performance in multi-hop retrieval tasks.
Knowledge Graph for Retrieval-Augmented Generation. Recently, Knowledge Graphs (KGs) have been proposed to facilitate multi-hop retrieval by explicitly modeling complex semantic relationships within and across documents (Peng et al., 2024; Procko & Ochoa, 2024). Current KG-based Retrieval-Augmented Generation (KG-RAG) can be categorized into two main approaches: Graph search based methods and Graph Neural Network (GNN) based methods. Graph search based methods identify seed nodes and leverage heuristic graph search algorithms (such as breadth-first search, community detection, or PageRank) to traverse the KG and retrieve relevant information for multi-hop questions (Edge et al., 2024; Guo et al., 2024; Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025). While these methods are intuitive, they are often susceptible to noise due to their dependence on heuristic strategies, which can introduce irrelevant nodes into the retrieval path. Additionally, the large search space of KGs amplifies the impact of noise, leading to information loss or inaccuracies. In contrast, GNN based methods utilize the message passing mechanism of GNNs to directly output retrieval results from the KG in an end-to-end manner (Luo et al., 2025; Wang et al., 2022) or to first identify seed nodes and then retrieve information using graph search algorithms (Mavromatis & Karypis, 2024; Fang et al., 2019; Yasunaga et al., 2021). These methods excel at modeling complex node dependencies through iterative information propagation; however, they also face challenges related to noise, as nodes may aggregate misleading information from irrelevant neighbors. Furthermore, GNN methods often struggle to capture multi-granular semantic information, limiting their effectiveness in addressing complex multi-hop questions.
3 Approach
This section presents the architecture of our proposed method, whose framework is shown in Figure  2. We first introduce the construction of the Multi-information Level Knowledge Graph (Multi-L KG). Next, we describe the detailed design of the Query-Specific Graph Neural Network (QSGNN). Finally, we present the pre-training strategy designed to enhance the performance of QSGNN in multi-hop question retrieval tasks.
Figure 2: Framework overview. It first constructs Multi-L KG to model the multi-level relationships within corpora. QSGNN is designed to aggregate information from different levels, all the aggregations are guided by query. After pre-training and fine-tuning, it can generate representations for multi-hop questions.
3.1 Multi-information level KG Construction
Multi-hop questions are inherently more semantically complex than one-hop questions, as they often encompass not only simple entities but also complex expressions whose meanings may extend beyond the coverage of entities within KGs. Current approaches primarily focus on entities while neglect the multi-granular information, limiting their effectiveness in handling multi-hop questions. To address this limitation, we propose the Multi-information Level Knowledge Graph (Multi-L KG), which integrates various granularities of information to model relationships across different semantic perspectives.
We define the Multi-L KG as $G=\left(\mathcal{O},\mathcal{C},\mathcal{D},\mathcal{E}_{oo},\mathcal{E}_{oc},\mathcal{E}_{od},\mathcal{E}_{cc},\mathcal{E}_{cd}\right)$. It contains three types of node sets: entity set $\mathcal{O}$, chunk set $\mathcal{C}$, document set $\mathcal{D}$ and five types of edge sets: entity-entity $\mathcal{E}_{oo}$, entity-chunk $\mathcal{E}_{oc}$, entity-document $\mathcal{E}_{od}$, chunk-chunk $\mathcal{E}_{cc}$, chunk-document $\mathcal{E}_{cd}$. The construction of Multi-L KG involves two main steps:
Node Extraction. Given the document corpus $\mathcal{D}$, we utilize $OpenIE(.)$  (Angeli et al., 2015; Etzioni et al., 2008; Zhou et al., 2022) to extract chunks from each document and derive triples from these chunks [^2] The prompt for the OpenIE extraction is detailed in the Appendix  A.10. Note that the term “chunk” can have various meanings and in this paper we view sentence as chunk because we think in most cases sentence is enough to act as a meaningful reference for questions. The subjects and the objects within the triples are collected as entities. The extraction process can be formalized as:
$OpenIE(\mathcal{D})\rightarrow\{(d_{i},\mathcal{C}_{d_{i}},\mathcal{T}_{d_{i}})|d_{i}\in\mathcal{D}\}\rightarrow\mathcal{C}=\{\mathcal{C}_{d_{i}}|d_{i}\in\mathcal{D}\},\mathcal{O}=\{o_{i}|o_{i}\in\mathcal{T}_{d_{i}},d_{i}\in\mathcal{D}\},$ (1)
where $d_{i}$ represents the document $i$, $\mathcal{C}_{d_{i}}$ denotes the set of chunks extracted from $d_{i}$, $\mathcal{T}_{d_{i}}$ is the set of triples derived from $d_{i}$, $o_{i}$ refers to the subject or object within these triples $\mathcal{T}_{d_{i}}$.
Relation Construction. The entity-entity set $\mathcal{E}_{oo}$ is established based on all extracted triples $\{\mathcal{T}_{d_{i}}|d_{i}\in\mathcal{D}\}$, capturing fundamental semantic connections. The chunk-chunk set $\mathcal{E}_{cc}$ is constructed by linking adjacent chunks within a document, as their adjacency reflects a kind of logical coherent expression within a document. The last three edge sets entity-chunk $\mathcal{E}_{oc}$, entity-document $\mathcal{E}_{od}$ and chunk-document $\mathcal{E}_{cd}$ are constructed based on containment relationships.
We argue that Multi-L KG provides a comprehensive framework for modeling relationships from multiple perspectives. First, it captures both basic semantic relationships($\mathcal{E}_{oo}$) and logical coherence($\mathcal{E}_{cc}$). Second, $\mathcal{E}_{oc}$ and $\mathcal{E}_{od}$ represent local-to-global relationships, enabling precise understanding of entities in different contexts (e.g., “Apple” as a fruit or a company). Finally, $\mathcal{E}_{cd}$ serves as a high-level semantic bridge, connecting chunks to their corresponding documents for broader contextual understanding.
3.2 Query-Specific Graph Neural Network
After constructing the Multi-L KG, the next key challenge is how to effectively retrieve relevant information for multi-hop questions. Existing methods face limitations in their capacity to integrate diverse information levels, and are likely to introduce irrelevant noise due to the complexity of Multi-L KG. To address these challenges, we propose a novel Query-Specific Graph Neural Network (QSGNN). This approach adaptively aggregates information based on the query, effectively mitigating noise while incorporating insights from all information levels within the Multi-L KG, to generate comprehensive, query-specific representations for retrieval. QSGNN has two kinds of message passing process: intra-level message passing and inter-level message passing. These two processes are described in detail as follows:
Intra-level Message Passing. Intra-level message passing aggregates information within the same level(i.e. $\mathcal{E}_{oo}$ and $\mathcal{E}_{cc}$). For each information level, the process is defined as:
$\alpha_{i,j}=Sim\left(\mathbf{h}^{l-1}_{i}W^{q}_{\alpha},\mathbf{h}^{l-1}_{j}W^{k}_{\alpha}\right),$ (2)
$\beta_{i,j}=Sim\left(\mathbf{q}W^{q}_{\beta},(\mathbf{h}^{l-1}_{i}||\mathbf{h}^{l-1}_{j})W^{k}_{\beta}\right),$
$attn=Softmax\left(\left\{\alpha_{i,j}+\beta_{i,j}|j\in\mathcal{N}(i)\right\}\right),$
$msg_{i}=\sum\limits_{j\in\mathcal{N}(i)}attn_{i,j}\mathbf{h}^{l-1}_{j}W^{v},$
$\mathbf{h}^{l}_{i}=P\left(Norm\left(\mathbf{h}^{l-1}_{i}+msg_{i}\right)\right),$
where $\mathbf{h}^{l}_{i/j}\in R^{n}$ is the representation of node $i/j$ in layer $l$, initialized with text embeddings from an embedding model  [^3] Embeddings are compressed to dimension $n$ via linear transformations as information bottlenecks.. $Sim(.)$ calculates the cosine similarity. $\mathbf{q}\in R^{n}$ is the query embedding.||$|$ represents the vector concatenation. $W^{q}_{\alpha/\beta}\in R^{n\times n}$, $W^{k}_{\alpha}\in R^{n\times n}$, $W^{k}_{\beta}\in R^{2n\times n}$, $W^{v}\in R^{n\times n}$ are all learnable parameters. $Norm(.)$ is the normalization function, $P(.)$ is the 2-layers MLP with ReLU  (Glorot et al., 2011).
Intra-level message passing captures two types of relationships in the Multi-L KG. One is the basic semantic relationship encoded in $\mathcal{E}_{oo}$ and the other is the logical coherence relationship represented by $\mathcal{E}_{cc}$. The aggregation considers not only connectivity, but also semantic similarity $\alpha_{i,j}$ and query alignment $\beta_{i,j}$, which minimizes noise while ensuring query-aware aggregation.
Inter-level Message Passing. Inter-level message passing aggregates information across different levels of Multi-L KG (i.e., $\mathcal{E}_{oc}$, $\mathcal{E}_{od}$ and $\mathcal{E}_{cd}$). For each aggregation, the process is defined as:
$\mathbf{p}_{i,j}=\left((\mathbf{h}^{l-1}_{i})W^{t}||(\mathbf{h}^{l-1}_{j})W^{s}\right),$ (3)
$\gamma_{i,j}=Sim\left(\mathbf{q}W^{q}_{\gamma},\mathbf{p}_{i,j}W^{k}_{\gamma}\right),$
$attn=Softmax\left(\left\{\gamma_{i,j}|j\in\mathcal{N}(i)\right\}\right),$
$\mathbf{msg}_{i}=\sum\limits_{j\in\mathcal{N}(i)}attn_{i,j}\mathbf{h}^{l-1}_{j}W^{v},$
$\mathbf{h}^{l}_{i}=P\left(Norm\left(\mathbf{h}^{l-1}_{i}+\mathbf{msg}_{i}\right)\right),$
where $W^{t}$ and $W^{s}$ are used to project heterogeneous nodes into a shared representation space. $W^{q}_{\gamma}\in R^{n\times n}$, $W_{\gamma}^{k}\in R^{2n\times n}$ are learnable parameters. The other notations follow those in Equation 2.
The inter-level message passing fuses the local, global relationships from hierarchical levels into representations, which contains comprehensive understanding towards the query.
Based on these two processes, QSGNN calculates the representations as follows:
$\mathbf{H}^{l}_{intra}=IntraMQ(\mathbf{q},\mathcal{O},\mathcal{C},\mathcal{E}_{oo},\mathcal{E}_{cc},\mathbf{H}^{l-1}),$ (4)
$\mathbf{H}^{l}=InterMQ(\mathbf{q},\mathcal{O},\mathcal{C},\mathcal{C},\mathcal{E}_{oc},\mathcal{E}_{od},\mathcal{E}_{cd},\mathbf{H}^{l}_{intra}),$
where $IntraMQ(.)$ refers to the intra-level message passing defined in Equation  2 and $InterMQ(.)$ is the inter-level message passing defined in Equation  3.
3.3 Training Strategy for QSGNN
To enhance the representation learning ability of QSGNN, we first pre-train it on synthesized data and then we fine-tune it with human annotations. The pre-training data are the synthesized (question, document) pairs extracted from the same corpora used to construct the Multi-L KG. These pairs contain one-hop questions and two-hop questions, which are all generated by OpenIE  [^4] The prompts can be found in Appendix  A.11.
One-hop Question. We generate one-hop questions based on the triples extracted from documents. For each triple-document pair $\{(sbj,verb,obj),d\}$, the question is generated in form of $(?,verb,obj)$ or $(sbj,verb,?)$ where the answer is sbjsbj or objobj and the support document is $d$.
Two-hop Question. Two-hop questions are generated based on relation chains formed by shared entities across different documents. For instance, if document $d_{i}$ contains the triple $(sbj_{i},verb_{i},ent)$ and document $d_{j}$ contains the triple $(ent,verb_{j},obj_{j})$ where entent is the common entity, we form the relation chain as $(sbj_{i},verb_{i},ent,verb_{j},obj_{j})$, the corresponding question is generated in the form $(?,verb_{i},ent,verb_{j},obj_{j})$ or $(sub_{i},verb_{i},ent,verb_{j},?)$ where the answer is $sub_{i}$ or $obj_{j}$ and the support documents are $d_{i}$ and $d_{j}$.
The QSGNN is pre-trained on the synthesized data using the NT-Xent loss  (Chen et al., 2020b):
$\mathcal{L}_{\text{NT-Xent}}=-\frac{1}{M}\sum_{i=1}^{M}\log\frac{\exp(\text{sim}(\mathbf{q}_{i},\mathbf{h}_{i})/\tau)}{\sum_{j\in Neg(i)}\mathbf{1}_{[j\neq i]}\exp(\text{sim}(\mathbf{q}_{i},\mathbf{h}_{j})/\tau)},$ (5)
where $\mathbf{q}_{i}$ denotes the query embedding, $h_{i}$ is the representation of support document, $M$ is the batch size, $\tau$ is the temperature parameter and we set it as 1.0 in our implementation, $Neg(.)$ denotes the negative sampling and we employ a hard negative sampling strategy  (Schroff et al., 2015; Xu et al., 2022) with sampling number set as 30.
After pre-training, we fine-tune QSGNN on the downstream retrieval task. In this stage, the questions are collected by humans with answers and support documents. We fine-tune QSGNN with the same loss function as Equation  5.
3.4 QSGNN Retrieval and Generation
Given a query $\mathbf{q}$, the retrieval process of QSGNN is denoted as:
$score=Sim(\mathbf{q},\mathbf{H}^{l}),$ (6)
$\mathcal{R}=TopK(\mathcal{D},score),$
where $\mathbf{H}^{l}$ represents the representations calculated by Equation 4. $TopK(.)$ selects the documents with $K$-highest scorescore. The retrieval results are feed into LLM as context to generate response:
$response=LLM(query,\mathcal{R}).$ (7)
4 Experimental Results
4.1 Experimental Settings
Dataset. We evaluate the effectiveness of our proposed method on three multi-hop QA benchmarks: MuSiQue  (Trivedi et al., 2022a), 2WikiMultiHopQA(2Wiki)  (Ho et al., 2020) and HotpotQA  (Yang et al., 2018). For a fair comparison, we use the same evaluation set(1000 samples for each dataset) as prior works  (Gutiérrez et al., 2025; Jimenez Gutierrez et al., 2024; Luo et al., 2025). Additionally, we sample extra 1,000 questions from the original dataset for fine-tuning and 225 questions for testing, without overlap with the evaluation set. The basic statistics of these dataset are shown in Table  1. Details about data process are presented in Appendix  A.2.
Table 1: Dataset statistics.
MuSiQue 2Wiki HotpotQA
#Entity 118,021 53,153 86,147
#Chunk 57,887 23,023 39,830
#Document 15,803 7,403 9,811
Baselines. We compare our method (QSGNN) with four baseline categories: i) Naive Retriever: we use BM25 (Robertson & Walker, 1994) and Contriever (Izacard et al., 2021) together with IRCoT (Trivedi et al., 2022b) as the naive retrievers. ii) Text Embedding: we use NV-Embed-v2-7B  (Lee et al., 2024) and GTE-Qwen2-7B-Instruct (Li et al., 2023) to generate embeddings for retrieval. iii) Graph Search: we use GraphRAG  (Edge et al., 2024), RAPTOR (Sarthi et al., 2024), LightRAG (Guo et al., 2024) HippoRAG (Jimenez Gutierrez et al., 2024) and HippoRAG2 (Gutiérrez et al., 2025) as the graph search baselines. iv) GNN Retriever: we use GNN-RAG (Mavromatis & Karypis, 2024) and GFM-RAG  (Luo et al., 2025) as the GNN retriever baselines. The implementation details for all baselines can be found in Appendix  A.3.
Metrics. Follow prior works  (Gutiérrez et al., 2025; Jimenez Gutierrez et al., 2024; Luo et al., 2025), we evaluate retrieval performance using recall@2 and recall@5 (top-5 retrieved documents) and Question-Answering(QA) performance with exact match (EM) and F1 scores.
Implementation Details. Our QSGNN is implemented with 2 layers, where each layer has one intra-leverl and one inter-level message passing process. Information dimension is set as 128. All the training uses 2 NVIDIA A100(80G) GPUs. In the pre-training stage, we use 95% samples for training and 5% samples for checkpoint selection [^5] The checkpoint is selected based on the retrieval performance on the 5% samples using Recall@5.. The max epoch number is set as 5. We set learning rate as 1e-4 without weight decay. The checkpoint is saved every 2000 steps. In the fine-tuning stage, the max epoch number is set as 3. We set learning rate as 5e-4. The checkpoint is saved every 100 steps and we select checkpoint based on test set (225 samples) to evaluate the performance. We use NV-Embed-v2-7B  (Lee et al., 2024) as the embedding model and we use Llama-3.3-70B-Instruct  (AI@Meta., 2024) as the OpenIE model. We retrieve the top-5 documents for LLM(Llama-3.3-70B-Instruct and GPT-4o-mini  (OpenAI., 2024)) as context for QA task.
4.2 Results and Discussion
Retrieval Task. Table 2 summarizes the retrieval performance across baselines. We have following key observations: i) QSGNN outperforms all baseline methods, especially compared to other GNN based methods, it achieves +11.9%, +12.9% and +5.4% improvements on MuSiQue, 2Wiki and HotpotQA respectively and +9.8% average improvement, showing its effectiveness on multi-hop questions. ii) The classic retrieval algorithms (BM25/Contriever) and text embedding baselines have limited performance as they fail to capture document relationships which is crucial for answering multi-hop questions. iii) Although GNN based methods are designed to capture multi-hop relationships, their message passing mechanisms may inadvertently aggregate irrelevant information, which degrades retrieval effectiveness. iv) Graph search methods generally perform better than other baselines, but their performance is still constrained by inherent limitations. The heuristic exploration rules often overlook multi-level information and may favor specific path patterns, resulting in data bias. For instance, the performance of RAPTOR varies a lot since it leverages Gaussian Mixture Model to detect clusters in KGs, which is highly biased on data.
Table 2: Retrieval performance of the baselines. Rec@2/Rec@5 denotes Recall@2/Recall@5. Average means the average performance across all the datasets. We highlight the best results with bold and the second best results with under line. We do not report the result of LightRAG or GraphRAG because they do not retrieve documents.
MuSiQue 2Wiki HotpotQA Average
Method Rec@2 Rec@5 Rec@2 Rec@5 Rec@2 Rec@5 Rec@2 Rec@5
BM25 32.76 43.62 55.32 65.31 57.36 75.82 48.48 61.58
Contriever+IRCoT 34.92 46.27 46.6 57.53 56.21 74.37 45.91 59.39
GTE-Qwen2-7B 51.24 67.72 66.73 77.78 77.24 90.15 65.07 78.55
NV-Embed-v2 50.12 68.19 67.16 78.66 79.27 92.82 65.52 79.89
GNN-RAG 34.13 53.72 58.67 69.14 67.18 81.69 53.33 68.18
GFM-RAG 49.66 67.19 62.92 79.20 74.05 88.86 62.21 78.42
RAPTOR 47.68 69.55 66.45 82.67 77.83 93.19 63.99 81.80
HippoRAG 47.55 68.15 71.95 86.25 70.17 84.11 63.22 79.50
HippoRAG2 50.42 72.29 73.31 88.91 78.25 92.56 67.33 84.59
QSGNN(Ours) 53.88 75.23 74.02 89.47 78.42 93.67 68.77 86.12
Table 3: QA performance using Llama-3.3-70B-Instruct. No Retrieval means the LLM response without external knowledge. Average means the average performance across all the datasets. We highlight the best results with bold and the second best results with under line.
MuSiQue 2Wiki HotpotQA Average
Method EM F1 EM F1 EM F1 EM F1
No Retrieval 17.29 26.26 36.37 41.87 36.42 46.69 30.03 38.27
BM25 20.38 28.04 43.15 49.49 51.26 61.44 38.26 46.32
Contriever+IRCoT 24.29 31.62 38.02 40.90 49.89 60.30 37.4 44.27
GTE-Qwen2-7B 33.60 40.9 54.11 59.61 58.61 70.97 48.77 57.16
NV-Embed-v2 31.66 39.26 53.78 59.99 59.53 72.27 48.32 57.17
GNN-RAG 28.29 34.09 38.46 49.77 52.03 63.52 39.59 49.13
GFM-RAG 30.05 39.06 48.99 59.58 55.60 67.86 44.88 55.5
RAPTOR 32.46 40.30 48.57 62.15 60.21 73.25 47.08 58.57
LightRAG 6.18 13.66 6.75 22.50 12.82 26.25 8.58 20.80
GraphRAG 34.10 41.90 52.46 63.63 59.89 73.35 48.82 59.63
HippoRAG2 35.16 42.44 56.60 66.18 60.92 74.06 50.89 60.89
QSGNN(Ours) 36.65 44.93 57.02 66.83 60.23 74.44 51.3 62.07
QA Task. We conduct QA experiments across the methods and the results are shown in Table 3. QSGNN outperforms other baselines due to its accurate retrieval. No Retrieval achieves low EM and F1 scores, demonstrating the importance of external knowledge for the LLM’s response. Notably, LightRAG struggles to perform well on these datasets, which is consistent with other reports  (Gutiérrez et al., 2025; Luo et al., 2025). The main reason is that LightRAG is incompatible with multi-hop retrieval workflows as discussed in  (Luo et al., 2025). The performance of the other methods have the consistent observations with Table 2 because the retrieval accuracy directly determines the quality of response. The results of GPT-4o-mini are shown in Appendix  A.4.
Multi-hop Performance. In order to analyze the detailed performance on multi-hop questions, we evaluate the performance of the competitive baselines on different hop numbers. The results of MuSiQue are shown in Table 4 and Figure  1. Firstly, we find that all the methods exhibit performance decline as hop number increases, because of the exponentially growing search space and noise accumulation. Secondly, GNN based methods particularly underperform in high-hop scenarios due to: i) the message passing aggregates the information of irrelevant nodes, ii) they fail to consider multiple information levels. Finally, as for the QSGNN, it achieves competitive performance in 2-hop questions compared to other baselines and the improvements become noticeable as the hop number increases, where it gets +33.8%, +85.8% and +231% improvements in Recall@5, F1 and EM on 4-hop questions respectively. This advantage stems from QSGNN’s query-guided attention mechanism and multi-level message passing, which significantly reduce the impact of noise and fuse the multi-information levels comprehensively(see the ablation study). The result of 2Wiki can be found in Appendix  A.5.
Table 4: Performance of different hop numbers. The best result is shown in bold.
MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
NV-Embed-v2 76.86 70.71 39.67 51.09 37.98 9.24 41.67 32.27 3.07
GFM-RAG 78.22 69.40 32.78 53.78 34.63 7.02 40.35 30.25 1.45
RAPTOR 77.39 71.25 44.86 53.20 38.35 8.57 42.82 33.14 2.76
GraphRAG - - - 54.99 39.91 9.79 45.18 34.38 3.22
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.2 35.76 3.78
QSGNN(Ours) 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
Table 5: Retrieval efficiency on MuSiQue. Time means the average retrieval time of each query.
Method NV-Embed-v2 GFM-RAG RAPTOR HippoRAG2 QSGNN
Recall@5 68.19 67.19 69.55 72.29 75.23
Time(s) 0.036 0.086 0.376 0.317 0.118
Efficiency. We evaluate the efficiency of retrieval across different methods, which is summarized in Table 5. Graph search based methods (HippoRAG2 and RAPTOR) exhibit higher time cost because the graph search strategies are expensive. GNN based methods (QSGNN and GFM-RAG) cost less time since message passing is an efficient way to aggregate graph information  (Hamilton et al., 2017). QSGNN shows more retrieval time than GFM-RAG due to the complexity of Multi-L KG.
Ablation Study. To evaluate the influence of different components, we conduct ablation studies and report the results in Table 6. Our key findings are: i) The model’s performance drops significantly without entities, as they form the basic semantic relationships in Multi-L KGs. Removing entities causes substantial knowledge loss. ii) No intra-level message passing severely impacts performance because QSGNN loses the ability to directly capture Multi-L KG structures but have to be bridged by chunk nodes. iii) The absence of query alignment has a particularly negative effect on high-hop questions, where the influence of noise will be amplified as the hop number increases. iv) Removing chunk nodes shows relatively less impact, as document nodes can provide global contextual information to compensate for their absence. More results are shown in Appendix  A.6.
Table 6: Ablation study of QSGNN. “w/o q_attn” means no alignment with query. “w/o intra” means no intra-level message passing. “w/o entity” and “w/o chunk” mean no entity or chunk in Multi-L KG. No inter-level message passing and No document are not reported here since we need retrieve documents.
MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
NV-Embed-v2 76.86 70.71 39.67 51.09 37.98 9.24 41.67 32.27 3.07
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
QSGNN 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
QSGNN(w/o entity) 60.34 43.31 25.02 27.25 16.56 2.05 19.58 9.65 1.13
QSGNN(w/o chunk) 77.65 74.99 62.60 51.18 45.82 16.55 42.74 35.69 9.08
QSGNN(w/o q_attn) 76.26 71.51 53.74 48.02 40.54 9.65 40.24 30.29 6.56
QSGNN(w/o intra) 71.33 68.42 43.15 39.88 33.39 6.93 28.76 22.85 2.61
Other Discussion. For the page limitation, we discuss the hyperparameter influence of QSGNN in Appendix  A.7. We conduct the good case study in Appendix  A.8. The bad case study and the discussion of limitations for QSGNN can be found in Appendix  A.9.
5 Conclusion
In this paper, we propose a novel graph representation learning framework for multi-hop questions retrieval. We first design a Multi-information Level Knowledge Graph (Multi-L KG) to model the multi-granular relationships among documents. Then we introduce Query-Specific Graph Neural Network (QSGNN). The QSGNN aggregates information from multi-levels, together with query alignment, to ensure comprehensive query-specific representations while filtering out the noise. The pre-training strategy further enhance the ability of QSGNN. Extensive experiments shows the effectiveness of our method in multi-hop questions.
Ethics Statement. We affirm that our research adheres to the highest ethical standards in the conduct of scientific research. This work does not involve human subjects, animal testing, harmful insights, discrimination, sensitive personal data, or any legal compliance problem. All experiments were conducted using publicly available datasets or Large Language Model (LLM) tools, and we have taken steps to ensure the reproducibility and transparency of our results. We have considered potential societal impacts of our research and believe that the proposed methods and findings contribute positively to the field of LLMs and its applications.
Reproducibility Statement. To ensure the reproducibility of our work, we provide the following statement. All code implementations of our proposed method are provided by an anonymous link, and we will make them publicly open source upon the acceptance of this paper. Additionally, all the large language models (LLMs) used in this paper are open source, as detailed in prior work  (Gutiérrez et al., 2025; Jimenez Gutierrez et al., 2024; Luo et al., 2025). Regarding the data used in our experiments, we follow the standard protocols established by open source works  (Gutiérrez et al., 2025). All the data processing details are presented in Appendix A.2. For better reproducibility, we also include example data in our code repository, which will be released publicly after the acceptance of the paper.
Appendix A Appendix
This supplementary material provides additional details on the proposed method and experimental results that could not be included in the main manuscript due to page limitations. Specifically, this appendix is organized as follows.
  • Sec. A.1 provides the use of Large Language Models (LLMs).
  • Sec. A.2 provides more details on the dataset settings and data processing.
  • Sec. A.3 outlines the implementation of baseline models and their experimental settings.
  • Sec. A.4 presents the QA performance on GPT-4o-mini.
  • Sec. A.5 provides the multi-hop performance on 2WikimultihopQA(2Wiki) dataset.
  • Sec. A.6 provides detailed ablation study for our method.
  • Sec. A.7 presents hyperparameter study for our method.
  • Sec. A.8 discusses three good cases of our method compared to HippoRAG2
  • Sec. A.9 presents three bad cases of our method and discusses the limitation and possible solution as future work.
  • Sec. A.10 presents the prompts used for OpenIE.
  • Sec. A.11 presents the prompts used for generating pre-training data.
A.1 The use of Large Language Models
We present the statement regarding the usage of Large Language Models (LLMs) in our paper. We assert that there is no involvement of LLMs in the innovation, implementation, or paper writing of our method. However, we still use LLMs in two ways:
i) Our proposed method is about the Retrieval-augmented generation (RAG) for LLMs. Therefore, in this paper, we use LLMs as the QA performer and the OpenIE tool to evaluate the effectiveness of our method, similar to other LLM-related works  (Guo et al., 2024; Jimenez Gutierrez et al., 2024; Luo et al., 2025).
ii) We use LLMs to help us correct some grammatical errors in our paper writing.
A.2 Dataset Settings
Table 7: Dataset statistics
MuSiQue 2Wiki HotpotQA
#Entity 118,021 53,153 86,147
#Chunk 57,887 23,023 39,830
#Document 15,803 7,403 9,811
#Pre-train 1-hop QA 91,621 52,122 73,128
#Pre-train 2-hop QA 58,923 12,097 20,619
#Pre-train QA 150,544 64,219 93,747
#Fine-tune 2-hop QA 270 782 1000
#Fine-tune 3-hop QA 482 0 0
#Fine-tune 4-hop QA 248 218 0
#Fine-tune QA 1,000 1000 1000
#Test 2-hop QA 66 172 225
#Test 3-hop QA 104 0 0
#Test 4-hop QA 55 53 0
#Test QA 225 225 225
#Eval 2-hop QA 488 847 1000
#Eval 3-hop QA 334 0 0
#Eval 4-hop QA 178 153 0
#Eval QA 1,000 1000 1000
Our dataset is constructed based on the benchmark datasets used by prior works (Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025; Luo et al., 2025). Here we present the construction details.
Dataset Construction. The evaluation set is the same as HippoRAG2(Gutiérrez et al., 2025), which is the state-of-the-art baseline. And the corpora is expanded. Since our method need extra QA pairs for fine-tuning, we randomly sample extra samples from each original dataset and we make sure there is no overlap with the evaluation set. For the MuSiQue, the sample should have at least 60% supported documents contained in existing corpora. For the 2Wiki, the sample should have at least 60% overlap with the corpora. For the HotpotQA, the overlap ratio is 50%. Within these extra samples, we randomly select 1000 questions for fine-tuning and 225 questions for testing. The corpora of each dataset will be expanded to cover the new samples related documents(supported documents and distracted documents).
Pre-training Data Construction. We generate synthesized pre-training data for each dataset. For the MuSiQue the one-hop QA pairs are 91,621 and the two-hop QA pairs are 58,923. For the 2Wiki, the one-hop QA pairs are 52,122 and the two-hop QA pairs are 12,097. For the HotpotQA, the one-hop QA pairs are 73,128 and the two-hop QA pairs are 20,619. We combine and shuffle all these QA pairs of each dataset as pre-training data.
The statistics of these datasets are shown in Table  7.
A.3 Baseline Settings
Implementation. The BM25 retrieval algorithm is implemented as BM25S (Lù, 2024). The Contriever  (Izacard et al., 2021) is implemented with the official code acting as the retrieval server in IRCoT  (Trivedi et al., 2022b). As for the text embedding based model, we use the NV-Embed-v2-7B and GTE-Qwen2-7B-Instruct from Huggingface  (Wolf et al., 2019). For the all the GNN based methods and graph search based methods we use their official implementations.
Experimental Settings. For all the baselines, we conduct experiment under the official settings or prior works experimental suggestions. Specifically, their settings are described as follows:
i) For BM25, we retrieve 5 most relevant documents based on the query terms frequency appearing in each document.
ii) For Contriever+IRCoT, we follow the suggestion of HippoRAG2  (Gutiérrez et al., 2025) that we retrieve the top 10 documents for each step, the maximum step number for MuSiQue is set as 4 and the maximum step number for other datasets is set as 2.
iii) For text embedding based methods, embeddings are calculated for all the documents within our datasets, we retrieve the 5 most relevant documents as  (Jimenez Gutierrez et al., 2024).
iv) For GFM-RAG, we use the official model implementation where we pre-train and fine-tune the GNN on our corpora and all the hyperparameters are set as  (Luo et al., 2025) suggests. We retrieve the 5 most relevant documents.
v) For GNN-RAG, we use ReaRev  (Mavromatis & Karypis, 2022) as GNN reasoning, it samples dense subgraph by PageRank  (Andersen et al., 2006) on our datasets, all the settings are the same as the official settings  (Mavromatis & Karypis, 2024).
vi) For GraphRAG and LightRAG, the implementations are based on the official codes, and the hyperparameters are set as HippoRAG2  (Gutiérrez et al., 2025) suggests.
vii) For RAPTOR, HippoRAG and HippoRAG2, the settings are set the same as their official implementations.
For all the methods, we use LLM (Llama-3.3-70B-Instruct  (AI@Meta., 2024) and GPT-4o-mini  (OpenAI., 2024)) for QA task.
A.4 QA Performance on GPT-4o-mini
Table 8: QA performance on GPT-4o-mini. Average means the average performance across all the datasets. We highlight the best results with bold and the second best results with under line.
MuSiQue 2Wiki HotpotQA Average
Method EM F1 EM F1 EM F1 EM F1
No Retrieval 10.81 21.39 31.83 36.51 27.44 41.47 23.36 33.12
BM25 12.54 23.88 32.25 40.30 34.50 52.62 26.43 38.93
Contriever+IRCoT 13.90 25.28 29.44 34.66 30.17 45.55 24.50 35.16
GTE-Qwen2-7B 31.15 39.88 50.75 52.37 52.19 66.17 44.70 52.81
NV-Embed-v2 28.07 37.15 49.22 51.51 54.85 68.43 44.05 52.36
GNN-RAG 23.02 30.12 34.52 48.25 49.12 58.99 35.55 45.79
GFM-RAG 28.30 36.81 43.79 53.26 51.86 64.73 41.32 51.6
RAPTOR 32.93 41.04 45.43 56.24 55.19 68.16 44.52 55.15
GraphRAG 33.31 40.13 50.99 58.08 56.45 69.24 46.92 55.82
HippoRAG2 33.12 40.30 51.49 60.18 56.39 70.53 47.00 57.00
QSGNN(Ours) 34.27 41.78 53.52 61.63 56.04 69.73 47.94 57.71
QA Task on GPT-4o-mini. Table 8 presents QA performance using GPT-4o-mini. We do not report the performance of LightRAG as it is unsuitable for multi-hop QA tasks. The results show that QSGNN achieves competitive EM and F1 scores, demonstrating its compatibility with different LLM backbones. No Retrieval gets the lowest EM and F1 score due to the limited knowledge of GPT-4o-mini. Notably, all methods show reduced performance compared to Llama-3-70B-Instruct(see Table  3), primarily due to the intrinsic knowledge gap between GPT-4o-mini and Llama-3.3-70B-Instruct. Other observations align with the report of Table  3.
A.5 Multi-hop Performance on 2Wiki
Table 9: Performance of different hop numbers on 2Wiki. The best result is shown in bold.
2Wiki(Recall@5) 2Wiki(F1) 2Wiki(EM)
Method 2-hop 4-hop 2-hop 4-hop 2-hop 4-hop
NV-Embed-v2 81.40 63.52 63.21 42.19 59.28 23.35
GFM-RAG 83.37 56.09 64.36 33.11 54.44 18.82
RAPTOR 85.91 64.75 66.13 40.14 53.71 20.12
GraphRAG - - 67.99 39.52 57.24 25.97
HippoRAG2 93.15 65.44 70.50 42.26 61.53 29.32
QSGNN 92.10 74.88 69.72 50.82 60.04 40.28
The multi-hop performance on 2Wiki dataset is presented in Table  9. We find that QSGNN can achieve +14.4%, +20.2% and +37.3% on Recall@5, F1 and EM respectively. For the other baselines, we have the same observations as Table  4.
A.6 Ablation Study
Table 10: Ablation study of QSGNN. “chunk” means performing QA task with chunk. “w/o inter” means no inter-level message passing. “w/o doc” means no document in Multi-L KG. “w/o pt” means no pre-training. “w/o ft” means no fine-tuning.
MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
NV-Embed-v2 76.86 70.71 39.67 51.09 37.98 9.24 41.67 32.27 3.07
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
QSGNN 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
QSGNN(w/o entity) 60.34 43.31 25.02 27.25 16.56 2.05 19.58 9.65 1.13
QSGNN(w/o chunk) 77.65 74.99 62.60 51.18 45.82 16.55 42.74 35.69 9.08
QSGNN(w/o q_attn) 76.26 71.51 53.74 48.02 40.54 9.65 40.24 30.29 6.56
QSGNN(w/o intra) 71.33 68.42 43.15 39.88 33.39 6.93 28.76 22.85 2.61
QSGNN(chunk) - - - 40.18 37.55 13.26 30.25 29.41 8.71
QSGNN(w/o inter) - - - 40.27 35.98 4.24 28.36 19.19 2.72
QSGNN(w/o doc) - - - 36.54 32.49 8.01 26.19 26.57 5.02
QSGNN(w/o pt) 63.66 53.93 36.83 35.64 28.14 5.15 24.62 19.28 2.53
QSGNN(w/o ft) 77.19 74.69 62.57 51.08 46.42 17.12 43.58 34.23 10.48
We present the ablation study of QSGNN on Table  10. As for the chunk retrieval (QSGNN + chunk, w/o inter or w/o doc), we retrieve the top 10 most relevant chunks for QA and we only report the EM and F1 score since we have no ground truth for chunk retrieval. We find that QSGNN(chunk) does not perform as well as QSGNN. It can be attributed to two factors: i) QSGNN is not directly trained on chunk labels. ii) Document generally contains more information than chunk. No inter-level message passing will not update the representations thus have no difference with the embeddings baselines, whose performance decrease a lot when hop number increases. No document lacks global context information, so it performs worse than QSGNN(chunk). In addition, QSGNN(w/o ft) can achieve competitive performance in 2-hop questions but fail to perform as well as QSGNN, since we pre-train QSGNN only on 1,2-hop questions. QSGNN(w/o pt) struggles to perform well because it can not be fitted well on only small number of training pairs.
A.7 Hyperparameter Discussion
Here we discussion some of important hyperparameters, including the receptive field (QSGNN layer number), negative sampling number, pre-training data scale and information bottleneck dimension.
Receptive Field Discussion. We discuss the relationships between GNN layers and the multi-hop performance, the results are shown in Table  11. We find that 1-layer QSGNN (one intra-level + one inter-level) have limited performance and the gap between 2-layer QSGNN becomes bigger as the hop number increase, it is because 1-layer QSGNN can only aggregate 2-hop information. The receptive field limits the potential of QSGNN. The 2-layer QSGNN achieves the best performance among 2,3,4 hop questions because it can aggregate more useful information. However, as the layer number increases(4 layers), the performance will suffer a lot because of the over-smoothing problem of GNN  (Rusch et al., 2023; Chen et al., 2020a).
Table 11: The influence of GNN layer number.
MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
QSGNN(1 layer) 73.81 69.14 55.13 46.25 41.86 11.46 38.29 29.60 7.83
QSGNN(2 layers) 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
QSGNN(3 layers) 76.35 74.55 61.66 51.34 45.24 16.45 42.11 34.50 9.13
QSGNN(4 layers) 62.60 58.29 35.37 38.93 29.15 4.61 26.68 18.68 2.03
Negative Sampling Number. Table 12 shows the impact of negative sampling number on model’s performance. According to the results, we find that low sampling number (Neg_N = 1) has limited performance due to inadequate discrimination between target documents and distractions. High sampling number (Neg_N=50, 100) shows limited benefits, because excessive easy negatives fail to provide meaningful guidance for QSGNN training.
Table 12: The influence of negative sampling number.
MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
QSGNN(Neg_N=1) 68.67 60.27 50.82 43.72 38.25 5.18 34.87 27.44 2.10
QSGNN(Neg_N=10) 77.97 76.16 62.14 51.03 46.68 17.62 42.28 38.54 10.76
QSGNN(Neg_N=30) 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
QSGNN(Neg_N=50) 77.18 76.69 63.51 51.34 46.99 18.36 44.07 38.02 11.65
QSGNN(Neg_N=100) 77.09 76.16 63.85 52.95 46.37 18.22 43.16 37.52 11.53
Figure 3: Influence of pre-training scale on various model dimension.
Pre-training Scale and Information Dimension. We pre-train QSGNN on the sythesized QA pairs from 0 to 150k across information dimension from 32 to 512 respectively, the results are shown in Figure  3. We conduct experiments on the MuSiQue dataset. As for the pre-training scale, the results show that insufficient pre-training data leads to sub-optimal performance across all the dimensions. As the amount of pre-training data increases, the performance of QSGNN gets better, although the marginal improvement decreases. As for the information dimension, the results indicate that increasing the information dimension improves the performance of QSGNN, as larger dimensions enhance model expressiveness through increased model scale. This observation is consistent with the model scaling analysis presented in GFM-RAG  (Luo et al., 2025). We believe that further improvements can be achieved by combining a larger model with a substantial amount of pre-training data. In this paper, we set the dimension as 128 considering the balance between performance and computational limitation.
A.8 Good Case Studies
We present three typical comparative cases between QSGNN and HippoRAG2 to demonstrate QSGNN’s advantages. In each case, correct retrieval evidence is marked in green and incorrect evidence in red. Notably, QSGNN’s marks are based on query alignment, whereas HippoRAG2’s marks is based on PPR scores  (Gutiérrez et al., 2025).
Figure 4: Good case 1 for QSGNN.
Figure 5: Good case 2 for QSGNN.
Figure 6: Good case 3 for QSGNN.
Good Case 1. Figure  4 illustrates a scenario where QSGNN achieves recall@5 of 1.0 with correct answer generation, while HippoRAG2 fails. The 4-hop question requires understanding composite semantics: i) “the only group larger than” corresponds to “the world’s second largest recorded music company”, ii) “the headquarters of” refers to “based in Santa Monica, California” and iii) “the explorer reach the city” denotes “set foot in the area was the party of explorer Gaspar de Portola”. QSGNN successfully identifies the critical documents describing Sony Music Entertainment’s ranking, location and the city information, whereas HippoRAG2 erroneously retrieves information about Jive Records’ headquarters. This failure stems from HippoRAG2’s selection of seed nodes (“headquarter”, “located in”) without high-level contextual information. The case validates Multi-L KG’s semantic modeling capability and query alignment’s effectiveness in filtering misinformation.
Good Case 2. Figure  5 shows QSGNN achieving perfect recall@5 while retrieval of HippoRAG2 is inaccurate. For the 4-hop question about Andrew Deveaux’s birthplace, QSGNN integrates evidence across multiple documents and each document describes one key evidence for the question: i) biographical data (“born in South Carolina”), ii) state capital history (“Columbia became capital in 1786”), and iii) geographic relationships (“Forest Acres borders Columbia”). The successful retrieval is due to the query-alignment and comprehensive understanding of contextual information. However, HippoRAG2 incorrectly retrieves Savannah county information due to over-reliance on “county” and “located on” seed nodes.
Good Case 3. Figure  6 presents a good case of QSGNN where it retrieves correctly and gives the correct answer but HippoRAG2 can’t answer correctly because of the inaccurate retrieval. In this case, QSGNN successfully find the snippets for the “where the Yongle Emperor greeted the Karmapa”, “the capitol city of Yaxing Coach’s headquarters location” and the history information about “Yangzhou”, which is the combination results of Multi-L KG and query-specific aggregation. HippoRAG2 retrieves redundant Sino-Tibetan relations documents. The failure occurs because HippoRAG2 prioritizes high PPR scores for entities “Yongle Emperor” and “Karmapa” without considering their contextual relationships.
A.9 Bad Case Discussion and Limitations
We analyze three representative failure cases of QSGNN in document retrieval, each illustrating distinct limitations that inform our future research directions.
Figure 7: Bad case 1 for QSGNN.
Figure 8: Bad case 2 for QSGNN.
Figure 9: Bad case 3 for QSGNN.
Bad Case 1. Figure  7 demonstrates QSGNN’s failure to retrieve documents containing “Freikorps” due to a misspelled query terminology (“free crops” should be “Freikorps”, we changed the “free crops” to “Freikorps” then QSGNN could retrieve correctly). This spelling error caused QSGNN to prioritize non-critical entities like “democratic government” and “Germany”. We also check the pre-training data and fine-tuning data, there is no information about the associations between “free crops” and “Freikorps”, which exacerbates the model’s confusion.
Bad Case 2. Figure  8 reveals QSGNN’s limitation in processing specific term (“San Clemente”, “Fuser”, “Alberto”), which may not be well represented by QSGNN since they are absent from both pre-training and fine-tuning corpora. The model deflect to general terms like “province” and “county”, retrieving irrelevant documents.
Bad Case 3. Figure  9 shows a typical bad case for QSGNN. In this case the term like “Messi” and “Barcelona” dwarf the key evidence “Diego Maradona, who Messi can bring comparison to”. The sequential dependency between finding “Diego Maradona” and subsequent evidence (“June 1982 transfer record”) further complicates retrieval. It is difficult for query alignment to identify the correct information because huge amount of documents describing “Messi” and “Barcelona”, which is also related to query, overwhelm the key snippets of “Diego Maradona”.
Limitation Discussion and Future Work. Based on the analysis of the bad cases above, QSGNN is not a silver bullet for all multi-hop scenarios. It exhibits limitations when dealing with specific terminologies especially when they are not well presented in the corpora (see bad case 1, 2), and also struggles to deal the situation where the prevalent mis-related documents overwhelm the key evidence with sequential answer dependencies (see bad case 3). We think these limitations can be mitigated by two strategies: i) For the first issue, selecting seed nodes as domain-specific terminologies and sampling constrained subgraphs(depending on the layer of QSGNN, in our case the longest hop is 4), could prioritize relevant documents in retrieval, though this risks information loss. ii) For the second issue, integrating Chain-of-Thought(CoT) reasoning  (Chen et al., 2019; Lyu et al., 2023) through query decomposition and iterative retrieval may mitigate the problem. However, the first solution is a tricky strategy and even may compromise the performance since subgraph sampling will lead to information loss. The second solution introduces CoT into framework, which is beyond the design topic of QSGNN. We will leave the combination of these two possible solutions without compromising the performance as the future work of QSGNN.
A.10 Prompts for OpenIE
We present the prompts we used for OpenIE workflow. Figure  10 shows the prompts used for sentence extraction, Figure  11 shows the prompts for entity extraction, and Figure  12 shows the prompts for triple extraction. First, we extract sentences from each document. Then we identify core entities from extracted sentences. At last, these entities are used to formulate triples from the document.
Figure 10: Sentence extraction prompt for OpenIE.
Figure 11: Entity extraction prompt for OpenIE.
Figure 12: Triple extraction prompt for OpenIE.
A.11 Prompts for Synthesized Pre-training Data
The prompts used for generate pre-training data are shown in Figure  13 and Figure  14.
Figure 13: One hop question generation prompt for OpenIE.
Figure 14: Two hop question generation prompt for OpenIE.

### v2

@@
-Query-Specific GNN: A Comprehensive Graph Representation Learning Method for Retrieval Augmented Generation
-Abstract
+Question-Adaptive Graph Learning for Multi-hop Retrieval Augmented Generation
+Abstract.
 Retrieval-augmented generation (RAG) has demonstrated its ability to enhance Large Language Models (LLMs) by integrating external knowledge sources.
@@
 We first introduce a Multi-information Level Knowledge Graph (Multi-L KG) to model various information levels for a more comprehensive understanding of multi-hop questions.
-Based on this, we design a Query-Specific Graph Neural Network (QSGNN) for representation learning on the Multi-L KG.
-QSGNN employs intra/inter-level message passing mechanisms, and in each message passing the information aggregation is guided by the query, which not only facilitates multi-granular information aggregation but also significantly reduces the impact of noise.
-To enhance its ability to learn robust representations, we further propose two synthesized data generation strategies for pre-training the QSGNN.
+Based on this, we design a Question-Adaptive Graph Neural Network (Quest-GNN) for representation learning on the Multi-L KG.
+Quest-GNN employs intra/inter-level message passing mechanisms, and in each message passing the information aggregation is guided by the question, which not only facilitates multi-granular information aggregation but also significantly reduces the impact of noise.
+To enhance its ability to learn robust representations, we further propose two synthesized data generation strategies for pre-training the Quest-GNN.
 Extensive experimental results demonstrate the effectiveness of our framework in multi-hop scenarios, especially in high-hop questions the improvement can reach 33.8%.
 The code is available at: https://github.com/Jerry2398/QSGNN.
-1 Introduction
-Retrieval-Augmented Generation (RAG) has emerged as a powerful paradigm for enhancing the capabilities of Large Language Models (LLMs)  (Peng et al., 2024; Procko & Ochoa, 2024; Gao et al., 2023).
-By retrieving pertinent information from external knowledge sources and integrating it into the generation process, RAG enables LLMs to ground their responses in factual information, significantly reducing the propensity for hallucinations  (Fan et al., 2024).
-A challenging scenario in RAG is handling multi-hop questions, where the answer cannot be directly supported by a single document [^1] Following prior works  (Edge et al., 2024; Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025), we view document as the retrieval target in this paper.
+Retrieval-augmented generation; Graph neural networks; Multi-hop Question
+journalyear: 2026copyright: ccconference: Proceedings of the 49th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 20–24, 2026; Melbourne, VIC, Australiabooktitle: Proceedings of the 49th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’26), July 20–24, 2026, Melbourne, VIC, Australiadoi: 10.1145/3805712.3809693isbn: 979-8-4007-2599-9/2026/07ccs: Information systems Information retrieval query processingccs: Information systems Information retrieval
+1. Introduction
+Retrieval-Augmented Generation (RAG) has emerged as a powerful paradigm for enhancing the capabilities of Large Language Models (LLMs)  (Peng et al., 2024; Gao et al., 2023; Liu et al., 2025, 2023; Hao and Zhang, 2023; Hao et al., 2025; Zhang et al., 2023).
+By retrieving pertinent information from external knowledge sources and integrating it into the generation process, RAG enables LLMs to ground their outputs in factual evidence, significantly reducing the propensity for hallucinations  (Fan et al., 2024; Procko and Ochoa, 2024).
+Figure 1. Performance on multi-hop questions. Existing methods struggle to perform well as hop number increases while Quest-GNN can achieve better performance on high-hop questions.
+A challenging scenario in RAG is handling multi-hop questions, where the answer cannot be directly supported by a single document [^1] Following prior works  (Edge et al., 2024; Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025), we view document as the retrieval unit in this paper.
 but requires information synthesized from multiple interrelated documents, each contributing a unique piece of the puzzle for the answer  (Dua et al., 2019; Chen et al., 2019).
 These questions demand that the RAG system accurately identify multiple relevant documents and synthesize them to form a coherent answer.
-Traditional RAG approaches, which primarily rely on semantic similarity for retrieval  (Karpukhin et al., 2020; Chen et al., 2024; Jiang et al., 2023; Trivedi et al., 2022b), often struggle with multi-hop questions since they overlook the relational dependencies between documents.
-To overcome this limitation, recent efforts have proposed to leverage Knowledge Graphs (KGs).
+Traditional RAG approaches, which tackle multi-hop questions by decomposing them into a sequence of simpler sub-questions for iterative retrieval (Trivedi et al., 2022b; Fang et al., 2025; Wang et al., 2025; Asai et al., 2024; Yu et al., 2024), often achieve suboptimal performance since they overlook the relational dependencies between documents and may suffer from accumulative intermediate inaccuracies (Hu and Lu, 2024; Arslan et al., 2024).
+To overcome this limitations, recent efforts have proposed to leverage Knowledge Graphs (KGs).
 These methods first construct KGs to model the relationships between documents.
-During retrieval stage, they either navigate the KGs using graph search strategies  (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2025) or utilize Graph Neural Networks (GNNs)  (Fang et al., 2019; Mavromatis & Karypis, 2024; Luo et al., 2025) to identify relevant information.
+During retrieval stage, they either navigate the KGs using graph search strategies  (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2025; Chen et al., 2025) or utilize Graph Neural Networks (GNNs)  (Fang et al., 2019; Mavromatis and Karypis, 2024; Luo et al., 2025; Yan et al., 2024; Zhang et al., 2024b; Hu et al., 2025) to identify relevant information.
 By using KGs, these approaches can better capture the dependencies between documents, thereby improving the retrieval process for multi-hop questions.
-Figure 1: Performance on multi-hop questions. Existing methods struggle to perform well as hop number increases while QSGNN can achieve better performance on high-hop questions.
 Although KG based methods have shown improved performance over traditional approaches, they still face significant challenges in addressing the complexities of multi-hop questions.
@@
 ii) Noise Sensitivity.
-The retrieval process for multi-hop questions is highly sensitive to noise since it has multiple retrieval targets.
+The retrieval process for multi-hop question is highly sensitive to noise since it has multiple retrieval targets.
 The noise in any one retrieval target will disrupt the final result.
-However, existing methods are mostly prone to noise influence.
+Existing methods are mostly prone to noise.
 For instance, graph search based methods rely on heuristic strategies to explore relevant information, which is likely to incorporate irrelevant nodes or miss important information.
@@
 This design encompasses the fine-grained information relationships and facilitates a more comprehensive understanding for multi-hop questions.
-ii) To learn representations from the Multi-L KG for multi-hop retrieval, we design a novel GNN model called Query-Specific Graph Neural Network (QSGNN).
+ii) To learn representations from the Multi-L KG for multi-hop retrieval, we design a novel GNN model called Question-Adaptive Graph Neural Network (Quest-GNN).
 Specifically, it adopts two kinds of message passing to handle the multiple information components of Multi-L KG: intra-level and inter-level message passing.
 Intra-level message passing focuses on the basic semantic relationships and the logical coherence relationships within each level, while inter-level message passing considers local-global relationships across levels.
-In each type of message passing, it considers both semantic relationships and alignment with the query.
+In each type of message passing, it considers both semantic relationships and alignment with the question.
 This design integrates multi-level information and significantly reduces the influence of irrelevant nodes, resulting in high-quality representations for multi-hop retrieval.
-iii) To further enhance the capability of QSGNN, we propose two data generation strategies to produce synthesized QA pairs for pre-training.
-This process requires no extra human cost and can be conducted after the construction of Multi-L KG.
-After pre-training, QSGNN can be quickly adapted to downstream tasks.
-We conduct extensive experiments to evaluate our method, which demonstrate that QSGNN achieves state-of-the-art performance compared to existing methods, especially in high hop questions the improvement can reach 33.8%.
-2 Related Work
-Retrieval-augmented generation(RAG).
-RAG aims at enhancing large language models (LLMs) by retrieving relevant information from external knowledge sources and integrating it into the generation process to enable fact-grounded responses  (Gao et al., 2023; Fan et al., 2024).
-Traditional RAG methods  (Chen et al., 2024; Li et al., 2023) encode documents from external knowledge sources into dense embeddings and retrieve relevant information by calculating similarity between query and document embeddings.
-However, these approaches fail to capture the complex relationships between documents, struggling to perform well on the multi-hop questions.
-In order to adapt the RAG to multi-hop questions, subsequent works  (Jiang et al., 2023; Trivedi et al., 2022b; Su et al., 2024) have proposed multi-step retrieval strategies, where they decompose the original question into multiple sub-questions via LLMs and retrieve relevant information for each sub-question.
-Although these methods explicitly model the internal structure of the query, they still neglect the multi-level relationships between documents, resulting in limited performance in multi-hop retrieval tasks.
-Knowledge Graph for Retrieval-Augmented Generation.
-Recently, Knowledge Graphs (KGs) have been proposed to facilitate multi-hop retrieval by explicitly modeling complex semantic relationships within and across documents (Peng et al., 2024; Procko & Ochoa, 2024).
+iii) To further enhance the Quest-GNN, we propose two data generation strategies to produce synthesized QA pairs for pre-training.
+This process requires no human cost and can be conducted after the construction of Multi-L KG.
+After pre-training, Quest-GNN can be quickly adapted to downstream tasks.
+We conduct extensive experiments to evaluate our method, which demonstrate that Quest-GNN achieves state-of-the-art performance compared to existing methods, especially in high hop questions the improvement can reach 33.8%.
+Figure 2.
+Framework overview.
+It first constructs Multi-L KG to model the multi-level relationships within corpora.
+Quest-GNN is designed to aggregate information from different levels, all the aggregations are guided by the question.
+After pre-training and fine-tuning, it can generate representations for multi-hop questions.
+2. Related Work
+Retrieval Augmented Generation (RAG).
+RAG aims at enhancing large language models (LLMs) by retrieving relevant information from external knowledge sources and integrating it into the generation process to enable fact-grounded responses  (Gao et al., 2023; Fan et al., 2024; Hao et al., 2023).
+Text embedding based methods  (Chen et al., 2024; Li et al., 2023; Liu, 2024) encode documents from external knowledge sources into dense embeddings and retrieve relevant information by calculating similarity between question and document embeddings.
+However, they fail to capture the complex relationships within the questions and the documents and thus struggle to perform well under multi-hop settings.
+In order to adapt the RAG to multi-hop questions, subsequent works  (Jiang et al., 2023; Trivedi et al., 2022b; Su et al., 2024; Fang et al., 2025; Wang et al., 2025; Asai et al., 2024; Yu et al., 2024) have proposed multi-step retrieval strategies, where they decompose the original question into multiple sub-questions via LLMs and retrieve relevant information for each sub-question iteratively.
+Although these methods explicitly model the internal structure of the question, they still neglect the multi-level relationships between documents, resulting in limited performance in multi-hop retrievals.
+Knowledge Graph for RAG.
+Recently, Knowledge Graphs (KGs) have been proposed to facilitate multi-hop retrieval by explicitly modeling complex semantic relationships within and across documents (Peng et al., 2024; Procko and Ochoa, 2024; Zhang et al., 2024a).
 Current KG-based Retrieval-Augmented Generation (KG-RAG) can be categorized into two main approaches: Graph search based methods and Graph Neural Network (GNN) based methods.
-Graph search based methods identify seed nodes and leverage heuristic graph search algorithms (such as breadth-first search, community detection, or PageRank) to traverse the KG and retrieve relevant information for multi-hop questions (Edge et al., 2024; Guo et al., 2024; Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025).
+Graph search based methods identify seed nodes and leverage heuristic graph search algorithms (such as breadth-first search, community detection, or LLM reasoning) to traverse the KG and retrieve relevant information for multi-hop questions (Edge et al., 2024; Guo et al., 2024; Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025; Sun et al., 2023; Ma et al., 2024; Luo et al., 2023).
 While these methods are intuitive, they are often susceptible to noise due to their dependence on heuristic strategies, which can introduce irrelevant nodes into the retrieval path.
 Additionally, the large search space of KGs amplifies the impact of noise, leading to information loss or inaccuracies.
-In contrast, GNN based methods utilize the message passing mechanism of GNNs to directly output retrieval results from the KG in an end-to-end manner (Luo et al., 2025; Wang et al., 2022) or to first identify seed nodes and then retrieve information using graph search algorithms (Mavromatis & Karypis, 2024; Fang et al., 2019; Yasunaga et al., 2021).
+In contrast, GNN based methods utilize the message passing mechanism of GNNs to directly output retrieval results from the KG in an end-to-end manner (Luo et al., 2025; Wang et al., 2022) or to first identify seed nodes and then retrieve information using graph search algorithms (Mavromatis and Karypis, 2024; Fang et al., 2019; Yasunaga et al., 2021; Li et al., 2025; He et al., 2024).
 These methods excel at modeling complex node dependencies through iterative information propagation; however, they also face challenges related to noise, as nodes may aggregate misleading information from irrelevant neighbors.
 Furthermore, GNN methods often struggle to capture multi-granular semantic information, limiting their effectiveness in addressing complex multi-hop questions.
-3 Approach
+3. Approach
 This section presents the architecture of our proposed method, whose framework is shown in Figure  2.
 We first introduce the construction of the Multi-information Level Knowledge Graph (Multi-L KG).
-Next, we describe the detailed design of the Query-Specific Graph Neural Network (QSGNN).
-Finally, we present the pre-training strategy designed to enhance the performance of QSGNN in multi-hop question retrieval tasks.
-Figure 2: Framework overview.
-It first constructs Multi-L KG to model the multi-level relationships within corpora.
-QSGNN is designed to aggregate information from different levels, all the aggregations are guided by query.
-After pre-training and fine-tuning, it can generate representations for multi-hop questions.
-3.1 Multi-information level KG Construction
+Next, we describe the detailed design of the Question-Adaptive Graph Neural Network (Quest-GNN).
+Finally, we present the pre-training strategy designed to enhance the performance of Quest-GNN in multi-hop question retrieval tasks.
+3.1. Multi-information level KG Construction
 Multi-hop questions are inherently more semantically complex than one-hop questions, as they often encompass not only simple entities but also complex expressions whose meanings may extend beyond the coverage of entities within KGs.
@@
 It contains three types of node sets: entity set $\mathcal{O}$, chunk set $\mathcal{C}$, document set $\mathcal{D}$ and five types of edge sets: entity-entity $\mathcal{E}_{oo}$, entity-chunk $\mathcal{E}_{oc}$, entity-document $\mathcal{E}_{od}$, chunk-chunk $\mathcal{E}_{cc}$, chunk-document $\mathcal{E}_{cd}$.
-The construction of Multi-L KG involves two main steps:
+The construction of Multi-L KG has two main steps:
 Node Extraction.
-Given the document corpus $\mathcal{D}$, we utilize $OpenIE(.)$  (Angeli et al., 2015; Etzioni et al., 2008; Zhou et al., 2022) to extract chunks from each document and derive triples from these chunks [^2] The prompt for the OpenIE extraction is detailed in the Appendix  A.10.
-Note that the term “chunk” can have various meanings and in this paper we view sentence as chunk because we think in most cases sentence is enough to act as a meaningful reference for questions.
+Given the document corpus $\mathcal{D}$, we utilize $OpenIE(.)$  (Angeli et al., 2015; Etzioni et al., 2008; Zhou et al., 2022) to extract chunks from each document and derive triples from these chunks [^2] The prompt for the OpenIE extraction is detailed in the Appendix  A.5.
+Note that the term “chunk” can have various meanings and in this paper we view sentence as chunk because generally sentence is enough to act as a meaningful reference for questions  (Guo et al., 2024).
 The subjects and the objects within the triples are collected as entities.
 The extraction process can be formalized as:
-$OpenIE(\mathcal{D})\rightarrow\{(d_{i},\mathcal{C}_{d_{i}},\mathcal{T}_{d_{i}})|d_{i}\in\mathcal{D}\}\rightarrow\mathcal{C}=\{\mathcal{C}_{d_{i}}|d_{i}\in\mathcal{D}\},\mathcal{O}=\{o_{i}|o_{i}\in\mathcal{T}_{d_{i}},d_{i}\in\mathcal{D}\},$ (1)
-where $d_{i}$ represents the document $i$, $\mathcal{C}_{d_{i}}$ denotes the set of chunks extracted from $d_{i}$, $\mathcal{T}_{d_{i}}$ is the set of triples derived from $d_{i}$, $o_{i}$ refers to the subject or object within these triples $\mathcal{T}_{d_{i}}$.
+(1) $OpenIE(\mathcal{D})\rightarrow\{(d_{i},\mathcal{C}_{d_{i}},\mathcal{T}_{d_{i}})|d_{i}\in\mathcal{D}\}$
+$\mathcal{C}=\{\mathcal{C}_{d_{i}}|d_{i}\in\mathcal{D}\},\mathcal{O}=\{o_{i}|o_{i}\in\mathcal{T}_{d_{i}},d_{i}\in\mathcal{D}\},$
+where $d_{i}$ represents the document $i$, $\mathcal{C}_{d_{i}}$ denotes the set of chunks extracted from $d_{i}$, $\mathcal{T}_{d_{i}}$ is the set of triples derived from $d_{i}$, $o_{i}$ refers to the subject or object within these triples $\mathcal{T}_{d_{i}}$, $\mathcal{C}$ denotes the chunk set extracted from $\mathcal{D}$ and $\mathcal{O}$ is the entity set.
 Relation Construction.
+There are five kinds of relationships ($\mathcal{E}_{oo}$, $\mathcal{E}_{cc}$, $\mathcal{E}_{oc}$, $\mathcal{E}_{od}$, $\mathcal{E}_{cd}$) need to be built.
 The entity-entity set $\mathcal{E}_{oo}$ is established based on all extracted triples $\{\mathcal{T}_{d_{i}}|d_{i}\in\mathcal{D}\}$, capturing fundamental semantic connections.
@@
 Finally, $\mathcal{E}_{cd}$ serves as a high-level semantic bridge, connecting chunks to their corresponding documents for broader contextual understanding.
-3.2 Query-Specific Graph Neural Network
+3.2. Question-Adaptive Graph Neural Network
 After constructing the Multi-L KG, the next key challenge is how to effectively retrieve relevant information for multi-hop questions.
 Existing methods face limitations in their capacity to integrate diverse information levels, and are likely to introduce irrelevant noise due to the complexity of Multi-L KG.
-To address these challenges, we propose a novel Query-Specific Graph Neural Network (QSGNN).
-This approach adaptively aggregates information based on the query, effectively mitigating noise while incorporating insights from all information levels within the Multi-L KG, to generate comprehensive, query-specific representations for retrieval.
-QSGNN has two kinds of message passing process: intra-level message passing and inter-level message passing.
+To address these challenges, we propose a novel Question-Adaptive Graph Neural Network (Quest-GNN).
+This approach adaptively aggregates information based on the question, effectively mitigating noise while incorporating insights from all information levels within the Multi-L KG, to generate comprehensive, question-adaptive representations for retrieval.
+Quest-GNN has two kinds of message passing process: intra-level message passing and inter-level message passing.
 These two processes are described in detail as follows:
@@
 For each information level, the process is defined as:
-$\alpha_{i,j}=Sim\left(\mathbf{h}^{l-1}_{i}W^{q}_{\alpha},\mathbf{h}^{l-1}_{j}W^{k}_{\alpha}\right),$ (2)
+(2) $\alpha_{i,j}=Sim\left(\mathbf{h}^{l-1}_{i}W^{q}_{\alpha},\mathbf{h}^{l-1}_{j}W^{k}_{\alpha}\right),$
 $\beta_{i,j}=Sim\left(\mathbf{q}W^{q}_{\beta},(\mathbf{h}^{l-1}_{i}||\mathbf{h}^{l-1}_{j})W^{k}_{\beta}\right),$
-$attn=Softmax\left(\left\{\alpha_{i,j}+\beta_{i,j}|j\in\mathcal{N}(i)\right\}\right),$
-$msg_{i}=\sum\limits_{j\in\mathcal{N}(i)}attn_{i,j}\mathbf{h}^{l-1}_{j}W^{v},$
+$intra\_attn=Softmax\left(\left\{\alpha_{i,j}+\beta_{i,j}|j\in\mathcal{N}(i)\right\}\right),$
+$msg_{i}=\sum\limits_{j\in\mathcal{N}(i)}intra\_attn_{i,j}\mathbf{h}^{l-1}_{j}W^{v},$
 $\mathbf{h}^{l}_{i}=P\left(Norm\left(\mathbf{h}^{l-1}_{i}+msg_{i}\right)\right),$
@@
 $Sim(.)$ calculates the cosine similarity.
-$\mathbf{q}\in R^{n}$ is the query embedding.||$|$ represents the vector concatenation.
+$\mathbf{q}\in R^{n}$ is the question embedding.||$|$ represents the vector concatenation.
 $W^{q}_{\alpha/\beta}\in R^{n\times n}$, $W^{k}_{\alpha}\in R^{n\times n}$, $W^{k}_{\beta}\in R^{2n\times n}$, $W^{v}\in R^{n\times n}$ are all learnable parameters.
@@
 One is the basic semantic relationship encoded in $\mathcal{E}_{oo}$ and the other is the logical coherence relationship represented by $\mathcal{E}_{cc}$.
-The aggregation considers not only connectivity, but also semantic similarity $\alpha_{i,j}$ and query alignment $\beta_{i,j}$, which minimizes noise while ensuring query-aware aggregation.
+The aggregation considers not only connectivity, but also semantic similarity $\alpha_{i,j}$ and question alignment $\beta_{i,j}$, which minimizes noise while ensuring question-aware aggregation.
 Inter-level Message Passing.
@@
 For each aggregation, the process is defined as:
-$\mathbf{p}_{i,j}=\left((\mathbf{h}^{l-1}_{i})W^{t}||(\mathbf{h}^{l-1}_{j})W^{s}\right),$ (3)
+(3) $\mathbf{p}_{i,j}=\left((\mathbf{h}^{l-1}_{i})W^{t}||(\mathbf{h}^{l-1}_{j})W^{s}\right),$
 $\gamma_{i,j}=Sim\left(\mathbf{q}W^{q}_{\gamma},\mathbf{p}_{i,j}W^{k}_{\gamma}\right),$
-$attn=Softmax\left(\left\{\gamma_{i,j}|j\in\mathcal{N}(i)\right\}\right),$
-$\mathbf{msg}_{i}=\sum\limits_{j\in\mathcal{N}(i)}attn_{i,j}\mathbf{h}^{l-1}_{j}W^{v},$
+$inter\_attn=Softmax\left(\left\{\gamma_{i,j}|j\in\mathcal{N}(i)\right\}\right),$
+$\mathbf{msg}_{i}=\sum\limits_{j\in\mathcal{N}(i)}inter\_attn_{i,j}\mathbf{h}^{l-1}_{j}W^{v},$
 $\mathbf{h}^{l}_{i}=P\left(Norm\left(\mathbf{h}^{l-1}_{i}+\mathbf{msg}_{i}\right)\right),$
@@
 The other notations follow those in Equation 2.
-The inter-level message passing fuses the local, global relationships from hierarchical levels into representations, which contains comprehensive understanding towards the query.
-Based on these two processes, QSGNN calculates the representations as follows:
-$\mathbf{H}^{l}_{intra}=IntraMQ(\mathbf{q},\mathcal{O},\mathcal{C},\mathcal{E}_{oo},\mathcal{E}_{cc},\mathbf{H}^{l-1}),$ (4)
-$\mathbf{H}^{l}=InterMQ(\mathbf{q},\mathcal{O},\mathcal{C},\mathcal{C},\mathcal{E}_{oc},\mathcal{E}_{od},\mathcal{E}_{cd},\mathbf{H}^{l}_{intra}),$
+The inter-level message passing fuses the local, global relationships from hierarchical levels into representations, which contains comprehensive understanding towards the question.
+Based on these two processes, Quest-GNN calculates the representations as follows:
+(4) $\mathbf{H}^{l}_{intra}=IntraMQ(\mathbf{q},\mathcal{O},\mathcal{C},\mathcal{E}_{oo},\mathcal{E}_{cc},\mathbf{H}^{l-1}),$
+$\mathbf{H}^{l}=InterMQ(\mathbf{q},\mathcal{O},\mathcal{C},\mathcal{D},\mathcal{E}_{oc},\mathcal{E}_{od},\mathcal{E}_{cd},\mathbf{H}^{l}_{intra}),$
 where $IntraMQ(.)$ refers to the intra-level message passing defined in Equation  2 and $InterMQ(.)$ is the inter-level message passing defined in Equation  3.
-3.3 Training Strategy for QSGNN
-To enhance the representation learning ability of QSGNN, we first pre-train it on synthesized data and then we fine-tune it with human annotations.
+3.3. Training Strategy for Quest-GNN
+To enhance the representation learning ability of Quest-GNN, we first pre-train it on synthesized data and then we fine-tune it with human annotations.
 The pre-training data are the synthesized (question, document) pairs extracted from the same corpora used to construct the Multi-L KG.
-These pairs contain one-hop questions and two-hop questions, which are all generated by OpenIE  [^4] The prompts can be found in Appendix  A.11.
+These pairs contain one-hop questions and two-hop questions, which are all generated by OpenIE  [^4] The prompts can be found in Appendix  A.5.
 One-hop Question.
 We generate one-hop questions based on the triples extracted from documents.
-For each triple-document pair $\{(sbj,verb,obj),d\}$, the question is generated in form of $(?,verb,obj)$ or $(sbj,verb,?)$ where the answer is sbjsbj or objobj and the support document is $d$.
+For each triple-document pair $\{(sbj,verb,obj),d\}$, the question is generated in form of $(?,verb,obj)$ or $(sbj,verb,?)$ where the answer is sbjsbj or objobj, the support document is $d$ and the ??
+denotes the question target asking the relation with $sbj/obj$.
 Two-hop Question.
 Two-hop questions are generated based on relation chains formed by shared entities across different documents.
-For instance, if document $d_{i}$ contains the triple $(sbj_{i},verb_{i},ent)$ and document $d_{j}$ contains the triple $(ent,verb_{j},obj_{j})$ where entent is the common entity, we form the relation chain as $(sbj_{i},verb_{i},ent,verb_{j},obj_{j})$, the corresponding question is generated in the form $(?,verb_{i},ent,verb_{j},obj_{j})$ or $(sub_{i},verb_{i},ent,verb_{j},?)$ where the answer is $sub_{i}$ or $obj_{j}$ and the support documents are $d_{i}$ and $d_{j}$.
-The QSGNN is pre-trained on the synthesized data using the NT-Xent loss  (Chen et al., 2020b):
-$\mathcal{L}_{\text{NT-Xent}}=-\frac{1}{M}\sum_{i=1}^{M}\log\frac{\exp(\text{sim}(\mathbf{q}_{i},\mathbf{h}_{i})/\tau)}{\sum_{j\in Neg(i)}\mathbf{1}_{[j\neq i]}\exp(\text{sim}(\mathbf{q}_{i},\mathbf{h}_{j})/\tau)},$ (5)
-where $\mathbf{q}_{i}$ denotes the query embedding, $h_{i}$ is the representation of support document, $M$ is the batch size, $\tau$ is the temperature parameter and we set it as 1.0 in our implementation, $Neg(.)$ denotes the negative sampling and we employ a hard negative sampling strategy  (Schroff et al., 2015; Xu et al., 2022) with sampling number set as 30.
-After pre-training, we fine-tune QSGNN on the downstream retrieval task.
+For instance, if document $d_{i}$ contains the triple $(sbj_{i},verb_{i},ent)$ and document $d_{j}$ contains the triple $(ent,verb_{j},obj_{j})$ where entent is the common entity, we form the chain as $(sbj_{i},verb_{i},ent,verb_{j},obj_{j})$, the corresponding question is generated from the relation chain in the form $(?,verb_{i},ent,verb_{j},obj_{j})$ or $(sbj_{i},verb_{i},ent,verb_{j},?)$ where the question is about the two-hop relationship between $sbj_{i}$ and $obj_{j}$, the support documents are $d_{i}$ and $d_{j}$ and the answer is $sbj_{i}$ or $obj_{j}$.
+The Quest-GNN is pre-trained on the synthesized data using the NT-Xent loss  (Chen et al., 2020b):
+(5) $\mathcal{L}_{\text{NT-Xent}}=-\frac{1}{M}\sum_{i=1}^{M}\log\frac{\exp(\text{sim}(\mathbf{q}_{i},\mathbf{h}_{i})/\tau)}{\sum_{j\in Neg(i)}\mathbf{1}_{[j\neq i]}\exp(\text{sim}(\mathbf{q}_{i},\mathbf{h}_{j})/\tau)},$
+where $\mathbf{q}_{i}$ denotes the question embedding, $h_{i}$ is the representation of support document, $M$ is the batch size, $\tau$ is the temperature parameter and we set it as 1.0 in our implementation, $Neg(.)$ denotes the negative sampling and we employ a hard negative sampling strategy  (Schroff et al., 2015; Xu et al., 2022) with sampling number set as 30.
+After pre-training, we fine-tune Quest-GNN on the downstream retrieval task.
 In this stage, the questions are collected by humans with answers and support documents.
-We fine-tune QSGNN with the same loss function as Equation  5.
-3.4 QSGNN Retrieval and Generation
-Given a query $\mathbf{q}$, the retrieval process of QSGNN is denoted as:
-$score=Sim(\mathbf{q},\mathbf{H}^{l}),$ (6)
+We fine-tune Quest-GNN with the same loss function as Equation  5.
+3.4. Quest-GNN Retrieval and Generation
+Given a question $\mathbf{q}$, the retrieval process of Quest-GNN is denoted as:
+(6) $score=Sim(\mathbf{q},\mathbf{H}^{l}),$
 $\mathcal{R}=TopK(\mathcal{D},score),$
@@
 The retrieval results are feed into LLM as context to generate response:
-$response=LLM(query,\mathcal{R}).$ (7)
-4 Experimental Results
-4.1 Experimental Settings
-Dataset.
+(7) $response=LLM(question,\mathcal{R}).$
+4. Experimental Results
+4.1. Experimental Settings
+4.1.1. Dataset
 We evaluate the effectiveness of our proposed method on three multi-hop QA benchmarks: MuSiQue  (Trivedi et al., 2022a), 2WikiMultiHopQA(2Wiki)  (Ho et al., 2020) and HotpotQA  (Yang et al., 2018).
 For a fair comparison, we use the same evaluation set (1000 samples for each dataset) as prior works  (Gutiérrez et al., 2025; Jimenez Gutierrez et al., 2024; Luo et al., 2025).
-Additionally, we sample extra 1,000 questions from the original dataset for fine-tuning and 225 questions for testing, without overlap with the evaluation set.
-The basic statistics of these dataset are shown in Table  1.
-Details about data process are presented in Appendix  A.2.
-Table 1: Dataset statistics.
+Additionally, we sample extra 1,000 questions from the original dataset for fine-tuning and 225 questions for validation, without overlap with the evaluation set.
+The basic statistics of these dataset are shown in Table  1 and detailed statistics and sampling strategy are presented in Appendix  A.1.
+Table 1. Dataset statistics.
 MuSiQue 2Wiki HotpotQA
@@
 #Document 15,803 7,403 9,811
-Baselines.
-We compare our method (QSGNN) with four baseline categories: i) Naive Retriever: we use BM25 (Robertson & Walker, 1994) and Contriever (Izacard et al., 2021) together with IRCoT (Trivedi et al., 2022b) as the naive retrievers.
-ii) Text Embedding: we use NV-Embed-v2-7B  (Lee et al., 2024) and GTE-Qwen2-7B-Instruct (Li et al., 2023) to generate embeddings for retrieval.
-iii) Graph Search: we use GraphRAG  (Edge et al., 2024), RAPTOR (Sarthi et al., 2024), LightRAG (Guo et al., 2024) HippoRAG (Jimenez Gutierrez et al., 2024) and HippoRAG2 (Gutiérrez et al., 2025) as the graph search baselines.
-iv) GNN Retriever: we use GNN-RAG (Mavromatis & Karypis, 2024) and GFM-RAG  (Luo et al., 2025) as the GNN retriever baselines.
-The implementation details for all baselines can be found in Appendix  A.3.
-Metrics.
+Table 2.
+Retrieval and QA performance using Llama-3.3-70B-Instruct.
+Rec@2/Rec@5 denotes Recall@2/Recall@5.
+No Retrieval means the LLM response without external knowledge.
+We highlight the best results with bold and the second best results with under line.
+We do not report the recall of LightRAG or GraphRAG because they do not retrieve documents.
+MuSiQue 2Wiki HotpotQA
+Method Rec@2 Rec@5 EM F1 Rec@2 Rec@5 EM F1 Rec@2 Rec@5 EM F1
+No Retrieval - - 17.29 26.26 - - 36.37 41.87 - - 36.42 46.69
+BM25 32.76 43.62 20.38 28.04 55.32 65.31 43.15 49.49 57.36 75.82 51.26 61.44
+Contriever+IRCoT 34.92 46.27 24.29 31.62 46.6 57.53 38.02 40.90 56.21 74.37 49.89 60.30
+GTE-Qwen2-7B 51.24 67.72 33.60 40.9 66.73 77.78 54.11 59.61 77.24 90.15 58.61 70.97
+NV-Embed-v2 50.12 68.19 31.66 39.26 67.16 78.66 53.78 59.99 79.27 92.82 59.53 72.27
+Auto-RAG 41.37 61.40 27.36 34.89 55.63 74.94 45.28 55.41 65.81 79.50 51.21 60.33
+Self-RAG 38.42 59.17 25.83 34.75 52.92 72.45 44.76 53.27 66.36 79.27 51.45 60.02
+KiRAG 42.96 63.05 28.66 36.35 58.67 79.01 48.91 59.53 67.18 81.69 52.03 63.52
+GNN-RAG 34.13 53.72 28.29 34.09 58.67 69.14 38.46 49.77 67.18 81.69 52.03 63.52
+G-retriever 47.14 66.46 30.13 38.04 62.38 78.05 46.84 58.10 74.16 89.49 54.73 65.90
+GFM-RAG 49.66 67.19 30.05 39.06 62.92 79.20 48.99 59.58 74.05 88.86 55.60 67.86
+GraphRAG - - 34.10 41.90 - - 52.46 63.63 - - 59.89 73.35
+LightRAG - - 6.18 13.66 - - 6.75 22.50 - - 12.82 26.25
+RAPTOR 47.68 69.55 32.46 40.30 66.45 82.67 48.57 62.15 77.83 93.19 60.21 73.25
+ToG2 46.16 68.04 32.75 40.28 64.86 81.22 47.74 62.14 75.14 90.18 57.99 72.53
+HippoRAG2 50.42 72.29 35.16 42.44 73.31 88.91 56.60 66.18 78.25 92.56 60.92 74.06
+Quest-GNN(Ours) 53.88 75.23 36.65 44.93 74.02 89.47 57.02 66.83 78.42 93.67 60.23 74.44
+4.1.2. Baselines
+We compare our method (Quest-GNN) with five baseline categories:
+  • Naive Retriever: we use BM25 (Robertson and Walker, 1994) and Contriever (Izacard et al., 2021) together with IRCoT (Trivedi et al., 2022b) as the naive retrievers.
+  • Text Embedding: we use NV-Embed-v2-7B  (Lee et al., 2024) and GTE-Qwen2-7B-Instruct (Li et al., 2023) to generate embeddings for retrieval.
+  • Iterative Retriever: we use Self-RAG  (Asai et al., 2024), Auto-RAG (Yu et al., 2024) and KiRAG  (Fang et al., 2025) as the multi-step RAG baselines.
+  • Graph Search: we use GraphRAG  (Edge et al., 2024), RAPTOR (Sarthi et al., 2024), LightRAG (Guo et al., 2024), ToG2 (Ma et al., 2024) and HippoRAG2 (Gutiérrez et al., 2025) as the graph search baselines.
+  • GNN Retriever: we use GNN-RAG (Mavromatis and Karypis, 2024), G-retriever (He et al., 2024) and GFM-RAG  (Luo et al., 2025) as the GNN retriever baselines.
+The settings of baselines can be found in Appendix  A.2.
+4.1.3. Metrics
 Follow prior works  (Gutiérrez et al., 2025; Jimenez Gutierrez et al., 2024; Luo et al., 2025), we evaluate retrieval performance using recall@2 and recall@5 (top-5 retrieved documents) and Question-Answering(QA) performance with exact match (EM) and F1 scores.
-Implementation Details.
-Our QSGNN is implemented with 2 layers, where each layer has one intra-leverl and one inter-level message passing process.
+4.1.4. Implementation Details
+Our Quest-GNN is implemented with 2 layers, where each layer has one intra-leverl and one inter-level message passing process.
 Information dimension is set as 128.
 All the training uses 2 NVIDIA A100(80G) GPUs.
-In the pre-training stage, we use 95% samples for training and 5% samples for checkpoint selection [^5] The checkpoint is selected based on the retrieval performance on the 5% samples using Recall@5..
+In the pre-training stage, we use 95% samples for training and 5% samples for checkpoint selection [^5] The checkpoint is selected based on the retrieval on the 5% samples using Recall@5..
 The max epoch number is set as 5.
@@
 We set learning rate as 5e-4.
-The checkpoint is saved every 100 steps and we select checkpoint based on test set (225 samples) to evaluate the performance.
+The checkpoint is saved every 100 steps and we select checkpoint based on dev set (225 samples) to evaluate the performance.
 We use NV-Embed-v2-7B  (Lee et al., 2024) as the embedding model and we use Llama-3.3-70B-Instruct  (AI@Meta., 2024) as the OpenIE model.
 We retrieve the top-5 documents for LLM(Llama-3.3-70B-Instruct and GPT-4o-mini  (OpenAI., 2024)) as context for QA task.
-4.2 Results and Discussion
-Retrieval Task.
-Table 2 summarizes the retrieval performance across baselines.
-We have following key observations: i) QSGNN outperforms all baseline methods, especially compared to other GNN based methods, it achieves +11.9%, +12.9% and +5.4% improvements on MuSiQue, 2Wiki and HotpotQA respectively and +9.8% average improvement, showing its effectiveness on multi-hop questions.
-ii) The classic retrieval algorithms (BM25/Contriever) and text embedding baselines have limited performance as they fail to capture document relationships which is crucial for answering multi-hop questions.
-iii) Although GNN based methods are designed to capture multi-hop relationships, their message passing mechanisms may inadvertently aggregate irrelevant information, which degrades retrieval effectiveness.
-iv) Graph search methods generally perform better than other baselines, but their performance is still constrained by inherent limitations.
+Table 3. Performance under different hop numbers. The best result is shown in bold.
+MuSiQue(Rec@5) MuSiQue(F1) MuSiQue(EM) 2Wiki(Rec@5) 2Wiki(F1) 2Wiki(EM)
+Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 4-hop 2-hop 4-hop 2-hop 4-hop
+NV-Embed-v2 76.86 70.71 39.67 51.09 37.98 9.24 41.67 32.27 3.07 81.40 63.52 63.21 42.19 59.28 23.35
+KiRAG 75.42 64.57 26.31 49.24 33.21 6.88 38.93 28.28 1.22 84.58 48.19 64.89 29.88 55.08 14.76
+G-retriever 79.16 67.41 29.86 54.29 31.06 6.60 41.02 29.38 1.66 82.36 54.19 63.05 30.67 52.48 15.63
+GFM-RAG 78.22 69.40 32.78 53.78 34.63 7.02 40.35 30.25 1.45 83.37 56.09 64.36 33.11 54.44 18.82
+GraphRAG - - - 54.99 39.91 9.79 45.18 34.38 3.22 - - 67.99 39.52 57.24 25.97
+RAPTOR 77.39 71.25 44.86 53.20 38.35 8.57 42.82 33.14 2.76 85.91 64.75 66.13 40.14 53.71 20.12
+ToG2 76.82 69.71 40.81 53.85 37.54 8.23 43.18 33.66 2.43 84.50 63.09 66.23 39.51 52.63 20.69
+HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.2 35.76 3.78 93.15 65.44 70.50 42.26 61.53 29.32
+Quest-GNN(Ours) 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50 92.10 74.88 69.72 50.82 60.04 40.28
+4.2. Retrieval and QA Performance
+Table 2 shows the retrieval and QA performance of all the methods. We have following observations:
+  • No Retrieval achieves low EM and F1 scores, demonstrating the importance of external knowledge for the LLM’s response.
+  • Quest-GNN outperforms all baseline methods, especially compared to other GNN based methods.
+Recall@5 achieves +11.9%, +12.9% and +5.4% improvements and F1 score gets +15.0%, +11.7% and +9.7% improvements on MuSiQue, 2Wiki and HotpotQA respectively, showing its effectiveness on multi-hop questions.
+  • The classic retrieval algorithms (BM25/Contriever) and text embedding baselines have limited performance as they fail to capture document relationships which is crucial for answering multi-hop questions.
+  • Iterative retrieval based methods show no benefit to multi-hop questions.
+On the one hand, they only model the question structure but ignore the multi-level relationships between documents.
+On the other hand, they significantly suffer from intermediate inaccuracies as the suboptimal results may deflect the whole reasoning process.
+  • Although GNN based methods are designed to capture multi-hop relationships, their message passing mechanisms may inadvertently aggregate irrelevant information, which degrades retrieval effectiveness and LLM’s response.
+  • Graph search based methods generally perform better than other baselines, but their performance is still constrained by inherent limitations.
 The heuristic exploration rules often overlook multi-level information and may favor specific path patterns, resulting in data bias.
 For instance, the performance of RAPTOR varies a lot since it leverages Gaussian Mixture Model to detect clusters in KGs, which is highly biased on data.
-Table 2: Retrieval performance of the baselines.
-Rec@2/Rec@5 denotes Recall@2/Recall@5.
-Average means the average performance across all the datasets.
-We highlight the best results with bold and the second best results with under line.
-We do not report the result of LightRAG or GraphRAG because they do not retrieve documents.
-MuSiQue 2Wiki HotpotQA Average
-Method Rec@2 Rec@5 Rec@2 Rec@5 Rec@2 Rec@5 Rec@2 Rec@5
-BM25 32.76 43.62 55.32 65.31 57.36 75.82 48.48 61.58
-Contriever+IRCoT 34.92 46.27 46.6 57.53 56.21 74.37 45.91 59.39
-GTE-Qwen2-7B 51.24 67.72 66.73 77.78 77.24 90.15 65.07 78.55
-NV-Embed-v2 50.12 68.19 67.16 78.66 79.27 92.82 65.52 79.89
-GNN-RAG 34.13 53.72 58.67 69.14 67.18 81.69 53.33 68.18
-GFM-RAG 49.66 67.19 62.92 79.20 74.05 88.86 62.21 78.42
-RAPTOR 47.68 69.55 66.45 82.67 77.83 93.19 63.99 81.80
-HippoRAG 47.55 68.15 71.95 86.25 70.17 84.11 63.22 79.50
-HippoRAG2 50.42 72.29 73.31 88.91 78.25 92.56 67.33 84.59
-QSGNN(Ours) 53.88 75.23 74.02 89.47 78.42 93.67 68.77 86.12
-Table 3: QA performance using Llama-3.3-70B-Instruct.
-No Retrieval means the LLM response without external knowledge.
-Average means the average performance across all the datasets.
-We highlight the best results with bold and the second best results with under line.
-MuSiQue 2Wiki HotpotQA Average
-Method EM F1 EM F1 EM F1 EM F1
-No Retrieval 17.29 26.26 36.37 41.87 36.42 46.69 30.03 38.27
-BM25 20.38 28.04 43.15 49.49 51.26 61.44 38.26 46.32
-Contriever+IRCoT 24.29 31.62 38.02 40.90 49.89 60.30 37.4 44.27
-GTE-Qwen2-7B 33.60 40.9 54.11 59.61 58.61 70.97 48.77 57.16
-NV-Embed-v2 31.66 39.26 53.78 59.99 59.53 72.27 48.32 57.17
-GNN-RAG 28.29 34.09 38.46 49.77 52.03 63.52 39.59 49.13
-GFM-RAG 30.05 39.06 48.99 59.58 55.60 67.86 44.88 55.5
-RAPTOR 32.46 40.30 48.57 62.15 60.21 73.25 47.08 58.57
-LightRAG 6.18 13.66 6.75 22.50 12.82 26.25 8.58 20.80
-GraphRAG 34.10 41.90 52.46 63.63 59.89 73.35 48.82 59.63
-HippoRAG2 35.16 42.44 56.60 66.18 60.92 74.06 50.89 60.89
-QSGNN(Ours) 36.65 44.93 57.02 66.83 60.23 74.44 51.3 62.07
-QA Task.
-We conduct QA experiments across the methods and the results are shown in Table 3.
-QSGNN outperforms other baselines due to its accurate retrieval.
-No Retrieval achieves low EM and F1 scores, demonstrating the importance of external knowledge for the LLM’s response.
 Notably, LightRAG struggles to perform well on these datasets, which is consistent with other reports  (Gutiérrez et al., 2025; Luo et al., 2025).
 The main reason is that LightRAG is incompatible with multi-hop retrieval workflows as discussed in  (Luo et al., 2025).
-The performance of the other methods have the consistent observations with Table 2 because the retrieval accuracy directly determines the quality of response.
-The results of GPT-4o-mini are shown in Appendix  A.4.
-Multi-hop Performance.
+We report the QA performance using GPT-4o-mini in Appendix  A.3.
+4.3. Multi-hop Performance
 In order to analyze the detailed performance on multi-hop questions, we evaluate the performance of the competitive baselines on different hop numbers.
-The results of MuSiQue are shown in Table 4 and Figure  1.
-Firstly, we find that all the methods exhibit performance decline as hop number increases, because of the exponentially growing search space and noise accumulation.
-Secondly, GNN based methods particularly underperform in high-hop scenarios due to: i) the message passing aggregates the information of irrelevant nodes, ii) they fail to consider multiple information levels.
-Finally, as for the QSGNN, it achieves competitive performance in 2-hop questions compared to other baselines and the improvements become noticeable as the hop number increases, where it gets +33.8%, +85.8% and +231% improvements in Recall@5, F1 and EM on 4-hop questions respectively.
-This advantage stems from QSGNN’s query-guided attention mechanism and multi-level message passing, which significantly reduce the impact of noise and fuse the multi-information levels comprehensively(see the ablation study).
-The result of 2Wiki can be found in Appendix  A.5.
-Table 4: Performance of different hop numbers. The best result is shown in bold.
-MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
-Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
-NV-Embed-v2 76.86 70.71 39.67 51.09 37.98 9.24 41.67 32.27 3.07
-GFM-RAG 78.22 69.40 32.78 53.78 34.63 7.02 40.35 30.25 1.45
-RAPTOR 77.39 71.25 44.86 53.20 38.35 8.57 42.82 33.14 2.76
-GraphRAG - - - 54.99 39.91 9.79 45.18 34.38 3.22
-HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.2 35.76 3.78
-QSGNN(Ours) 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
-Table 5: Retrieval efficiency on MuSiQue. Time means the average retrieval time of each query.
-Method NV-Embed-v2 GFM-RAG RAPTOR HippoRAG2 QSGNN
-Recall@5 68.19 67.19 69.55 72.29 75.23
-Time(s) 0.036 0.086 0.376 0.317 0.118
-Efficiency.
+The results are shown in Table 3 and we visualize the performance on MuSiQue as Figure  1.
+We make the following observations:
+  • All the methods exhibit performance decline as hop number increases, because of the exponentially growing search space and noise accumulation.
+  • Iterative retrieval based methods underperform in high-hop scenarios due to: i) the negelect of complex relationships between documents, ii) accumulated errors from intermediate results.
+  • Compared to graph search based methods, existing GNN based methods have more performance decrease in high-hop questions because the message passing aggregates the information of irrelevant nodes without question alignment.
+  • As for Quest-GNN, it achieves competitive performance in 2-hop questions compared to other baselines and the improvements become noticeable as the hop number increases.
+For example, Quest-GNN gets +33.8%, +85.8% and +231% improvements respectively in terms of Recall@5, F1 and EM on 4-hop questions of MuSiQue.
+This advantage stems from Quest-GNN’s question-guided attention mechanism and multi-level message passing, which significantly reduce the impact of noise and fuse the multi-information levels comprehensively.
+4.4. Efficiency
 We evaluate the efficiency of retrieval across different methods, which is summarized in Table 4.
+Iterative retrieval based methods (Auto-RAG and KiRAG) show high efficiency because they have no extra cost of indexing or searching on graph.
 Graph search based methods (HippoRAG2 and RAPTOR) exhibit higher time cost because the graph search strategies are expensive.
-GNN based methods (QSGNN and GFM-RAG) cost less time since message passing is an efficient way to aggregate graph information  (Hamilton et al., 2017).
-QSGNN shows more retrieval time than GFM-RAG due to the complexity of Multi-L KG.
+GNN based methods (Quest-GNN, G-retriever and GFM-RAG) cost less time since message passing is an efficient way to aggregate graph information  (Hamilton et al., 2017).
+Quest-GNN shows more retrieval time than GFM-RAG and G-retriever due to the complexity of Multi-L KG.
+Table 4. Retrieval efficiency on MuSiQue. Time means the average retrieval time of each question.
+Method Recall@5 Time(s)
+NV-Embed-v2 68.19 0.036
+Auto-RAG 61.40 0.056
+KiRAG 63.05 0.072
+G-retriever 66.46 0.051
+GFM-RAG 67.19 0.086
+RAPTOR 69.55 0.376
+HippoRAG2 72.29 0.317
+Quest-GNN 75.23 0.118
+Table 5.
 Ablation study.
-To evaluate the influence of different components, we conduct ablation studies and report the results in Table 6.
-Our key findings are: i) The model’s performance drops significantly without entities, as they form the basic semantic relationships in Multi-L KGs.
-Removing entities causes substantial knowledge loss.
-ii) No intra-level message passing severely impacts performance because QSGNN loses the ability to directly capture Multi-L KG structures but have to be bridged by chunk nodes.
-iii) The absence of query alignment has a particularly negative effect on high-hop questions, where the influence of noise will be amplified as the hop number increases.
-iv) Removing chunk nodes shows relatively less impact, as document nodes can provide global contextual information to compensate for their absence.
-More results are shown in Appendix  A.6.
-Table 6: Ablation study of QSGNN.
-“w/o q_attn” means no alignment with query.
-“w/o intra” means no intra-level message passing.
-“w/o entity” and “w/o chunk” mean no entity or chunk in Multi-L KG.
-No inter-level message passing and No document are not reported here since we need retrieve documents.
-MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
-Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
-NV-Embed-v2 76.86 70.71 39.67 51.09 37.98 9.24 41.67 32.27 3.07
-HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
-QSGNN 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
-QSGNN(w/o entity) 60.34 43.31 25.02 27.25 16.56 2.05 19.58 9.65 1.13
-QSGNN(w/o chunk) 77.65 74.99 62.60 51.18 45.82 16.55 42.74 35.69 9.08
-QSGNN(w/o q_attn) 76.26 71.51 53.74 48.02 40.54 9.65 40.24 30.29 6.56
-QSGNN(w/o intra) 71.33 68.42 43.15 39.88 33.39 6.93 28.76 22.85 2.61
-Other Discussion.
-For the page limitation, we discuss the hyperparameter influence of QSGNN in Appendix  A.7.
-We conduct the good case study in Appendix  A.8.
-The bad case study and the discussion of limitations for QSGNN can be found in Appendix  A.9.
-5 Conclusion
-In this paper, we propose a novel graph representation learning framework for multi-hop questions retrieval.
-We first design a Multi-information Level Knowledge Graph (Multi-L KG) to model the multi-granular relationships among documents.
-Then we introduce Query-Specific Graph Neural Network (QSGNN).
-The QSGNN aggregates information from multi-levels, together with query alignment, to ensure comprehensive query-specific representations while filtering out the noise.
-The pre-training strategy further enhance the ability of QSGNN.
-Extensive experiments shows the effectiveness of our method in multi-hop questions.
-Ethics Statement.
-We affirm that our research adheres to the highest ethical standards in the conduct of scientific research.
-This work does not involve human subjects, animal testing, harmful insights, discrimination, sensitive personal data, or any legal compliance problem.
-All experiments were conducted using publicly available datasets or Large Language Model (LLM) tools, and we have taken steps to ensure the reproducibility and transparency of our results.
-We have considered potential societal impacts of our research and believe that the proposed methods and findings contribute positively to the field of LLMs and its applications.
-Reproducibility Statement.
-To ensure the reproducibility of our work, we provide the following statement.
-All code implementations of our proposed method are provided by an anonymous link, and we will make them publicly open source upon the acceptance of this paper.
-Additionally, all the large language models (LLMs) used in this paper are open source, as detailed in prior work  (Gutiérrez et al., 2025; Jimenez Gutierrez et al., 2024; Luo et al., 2025).
-Regarding the data used in our experiments, we follow the standard protocols established by open source works  (Gutiérrez et al., 2025).
-All the data processing details are presented in Appendix A.2.
-For better reproducibility, we also include example data in our code repository, which will be released publicly after the acceptance of the paper.
-Appendix A Appendix
-This supplementary material provides additional details on the proposed method and experimental results that could not be included in the main manuscript due to page limitations.
-Specifically, this appendix is organized as follows.
-  • Sec. A.1 provides the use of Large Language Models (LLMs).
-  • Sec. A.2 provides more details on the dataset settings and data processing.
-  • Sec. A.3 outlines the implementation of baseline models and their experimental settings.
-  • Sec. A.4 presents the QA performance on GPT-4o-mini.
-  • Sec. A.5 provides the multi-hop performance on 2WikimultihopQA(2Wiki) dataset.
-  • Sec. A.6 provides detailed ablation study for our method.
-  • Sec. A.7 presents hyperparameter study for our method.
-  • Sec. A.8 discusses three good cases of our method compared to HippoRAG2
-  • Sec. A.9 presents three bad cases of our method and discusses the limitation and possible solution as future work.
-  • Sec. A.10 presents the prompts used for OpenIE.
-  • Sec. A.11 presents the prompts used for generating pre-training data.
-A.1 The use of Large Language Models
-We present the statement regarding the usage of Large Language Models (LLMs) in our paper.
-We assert that there is no involvement of LLMs in the innovation, implementation, or paper writing of our method.
-However, we still use LLMs in two ways:
-i) Our proposed method is about the Retrieval-augmented generation (RAG) for LLMs.
-Therefore, in this paper, we use LLMs as the QA performer and the OpenIE tool to evaluate the effectiveness of our method, similar to other LLM-related works  (Guo et al., 2024; Jimenez Gutierrez et al., 2024; Luo et al., 2025).
-ii) We use LLMs to help us correct some grammatical errors in our paper writing.
-A.2 Dataset Settings
-Table 7: Dataset statistics
-MuSiQue 2Wiki HotpotQA
-#Entity 118,021 53,153 86,147
-#Chunk 57,887 23,023 39,830
-#Document 15,803 7,403 9,811
-#Pre-train 1-hop QA 91,621 52,122 73,128
-#Pre-train 2-hop QA 58,923 12,097 20,619
-#Pre-train QA 150,544 64,219 93,747
-#Fine-tune 2-hop QA 270 782 1000
-#Fine-tune 3-hop QA 482 0 0
-#Fine-tune 4-hop QA 248 218 0
-#Fine-tune QA 1,000 1000 1000
-#Test 2-hop QA 66 172 225
-#Test 3-hop QA 104 0 0
-#Test 4-hop QA 55 53 0
-#Test QA 225 225 225
-#Eval 2-hop QA 488 847 1000
-#Eval 3-hop QA 334 0 0
-#Eval 4-hop QA 178 153 0
-#Eval QA 1,000 1000 1000
-Our dataset is constructed based on the benchmark datasets used by prior works (Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025; Luo et al., 2025). Here we present the construction details.
-Dataset Construction.
-The evaluation set is the same as HippoRAG2(Gutiérrez et al., 2025), which is the state-of-the-art baseline.
-And the corpora is expanded.
-Since our method need extra QA pairs for fine-tuning, we randomly sample extra samples from each original dataset and we make sure there is no overlap with the evaluation set.
-For the MuSiQue, the sample should have at least 60% supported documents contained in existing corpora.
-For the 2Wiki, the sample should have at least 60% overlap with the corpora.
-For the HotpotQA, the overlap ratio is 50%.
-Within these extra samples, we randomly select 1000 questions for fine-tuning and 225 questions for testing.
-The corpora of each dataset will be expanded to cover the new samples related documents(supported documents and distracted documents).
-Pre-training Data Construction.
-We generate synthesized pre-training data for each dataset.
-For the MuSiQue the one-hop QA pairs are 91,621 and the two-hop QA pairs are 58,923.
-For the 2Wiki, the one-hop QA pairs are 52,122 and the two-hop QA pairs are 12,097.
-For the HotpotQA, the one-hop QA pairs are 73,128 and the two-hop QA pairs are 20,619.
-We combine and shuffle all these QA pairs of each dataset as pre-training data.
-The statistics of these datasets are shown in Table  7.
-A.3 Baseline Settings
-Implementation.
-The BM25 retrieval algorithm is implemented as BM25S (Lù, 2024).
-The Contriever  (Izacard et al., 2021) is implemented with the official code acting as the retrieval server in IRCoT  (Trivedi et al., 2022b).
-As for the text embedding based model, we use the NV-Embed-v2-7B and GTE-Qwen2-7B-Instruct from Huggingface  (Wolf et al., 2019).
-For the all the GNN based methods and graph search based methods we use their official implementations.
-Experimental Settings. For all the baselines, we conduct experiment under the official settings or prior works experimental suggestions. Specifically, their settings are described as follows:
-i) For BM25, we retrieve 5 most relevant documents based on the query terms frequency appearing in each document.
-ii) For Contriever+IRCoT, we follow the suggestion of HippoRAG2  (Gutiérrez et al., 2025) that we retrieve the top 10 documents for each step, the maximum step number for MuSiQue is set as 4 and the maximum step number for other datasets is set as 2.
-iii) For text embedding based methods, embeddings are calculated for all the documents within our datasets, we retrieve the 5 most relevant documents as  (Jimenez Gutierrez et al., 2024).
-iv) For GFM-RAG, we use the official model implementation where we pre-train and fine-tune the GNN on our corpora and all the hyperparameters are set as  (Luo et al., 2025) suggests.
-We retrieve the 5 most relevant documents.
-v) For GNN-RAG, we use ReaRev  (Mavromatis & Karypis, 2022) as GNN reasoning, it samples dense subgraph by PageRank  (Andersen et al., 2006) on our datasets, all the settings are the same as the official settings  (Mavromatis & Karypis, 2024).
-vi) For GraphRAG and LightRAG, the implementations are based on the official codes, and the hyperparameters are set as HippoRAG2  (Gutiérrez et al., 2025) suggests.
-vii) For RAPTOR, HippoRAG and HippoRAG2, the settings are set the same as their official implementations.
-For all the methods, we use LLM (Llama-3.3-70B-Instruct  (AI@Meta., 2024) and GPT-4o-mini  (OpenAI., 2024)) for QA task.
-A.4 QA Performance on GPT-4o-mini
-Table 8: QA performance on GPT-4o-mini. Average means the average performance across all the datasets. We highlight the best results with bold and the second best results with under line.
-MuSiQue 2Wiki HotpotQA Average
-Method EM F1 EM F1 EM F1 EM F1
-No Retrieval 10.81 21.39 31.83 36.51 27.44 41.47 23.36 33.12
-BM25 12.54 23.88 32.25 40.30 34.50 52.62 26.43 38.93
-Contriever+IRCoT 13.90 25.28 29.44 34.66 30.17 45.55 24.50 35.16
-GTE-Qwen2-7B 31.15 39.88 50.75 52.37 52.19 66.17 44.70 52.81
-NV-Embed-v2 28.07 37.15 49.22 51.51 54.85 68.43 44.05 52.36
-GNN-RAG 23.02 30.12 34.52 48.25 49.12 58.99 35.55 45.79
-GFM-RAG 28.30 36.81 43.79 53.26 51.86 64.73 41.32 51.6
-RAPTOR 32.93 41.04 45.43 56.24 55.19 68.16 44.52 55.15
-GraphRAG 33.31 40.13 50.99 58.08 56.45 69.24 46.92 55.82
-HippoRAG2 33.12 40.30 51.49 60.18 56.39 70.53 47.00 57.00
-QSGNN(Ours) 34.27 41.78 53.52 61.63 56.04 69.73 47.94 57.71
-QA Task on GPT-4o-mini.
-Table 8 presents QA performance using GPT-4o-mini.
-We do not report the performance of LightRAG as it is unsuitable for multi-hop QA tasks.
-The results show that QSGNN achieves competitive EM and F1 scores, demonstrating its compatibility with different LLM backbones.
-No Retrieval gets the lowest EM and F1 score due to the limited knowledge of GPT-4o-mini.
-Notably, all methods show reduced performance compared to Llama-3-70B-Instruct(see Table  3), primarily due to the intrinsic knowledge gap between GPT-4o-mini and Llama-3.3-70B-Instruct.
-Other observations align with the report of Table  3.
-A.5 Multi-hop Performance on 2Wiki
-Table 9: Performance of different hop numbers on 2Wiki. The best result is shown in bold.
-2Wiki(Recall@5) 2Wiki(F1) 2Wiki(EM)
-Method 2-hop 4-hop 2-hop 4-hop 2-hop 4-hop
-NV-Embed-v2 81.40 63.52 63.21 42.19 59.28 23.35
-GFM-RAG 83.37 56.09 64.36 33.11 54.44 18.82
-RAPTOR 85.91 64.75 66.13 40.14 53.71 20.12
-GraphRAG - - 67.99 39.52 57.24 25.97
-HippoRAG2 93.15 65.44 70.50 42.26 61.53 29.32
-QSGNN 92.10 74.88 69.72 50.82 60.04 40.28
-The multi-hop performance on 2Wiki dataset is presented in Table  9.
-We find that QSGNN can achieve +14.4%, +20.2% and +37.3% on Recall@5, F1 and EM respectively.
-For the other baselines, we have the same observations as Table  4.
-A.6 Ablation Study
-Table 10: Ablation study of QSGNN.
 “chunk” means performing QA task with chunk.
-“w/o inter” means no inter-level message passing.
-“w/o doc” means no document in Multi-L KG.
+“w/o entity”, “w/o chunk” and “w/o doc” mean no entity, chunk or document in Multi-L KG.
+“w/o intra” and “w/o inter” mean no intra-level message passing or no inter-level message passing.
+“w/o q_attn” means no question alignment.
 “w/o pt” means no pre-training.
@@
 Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
-NV-Embed-v2 76.86 70.71 39.67 51.09 37.98 9.24 41.67 32.27 3.07
-HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
-QSGNN 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
-QSGNN(w/o entity) 60.34 43.31 25.02 27.25 16.56 2.05 19.58 9.65 1.13
-QSGNN(w/o chunk) 77.65 74.99 62.60 51.18 45.82 16.55 42.74 35.69 9.08
-QSGNN(w/o q_attn) 76.26 71.51 53.74 48.02 40.54 9.65 40.24 30.29 6.56
-QSGNN(w/o intra) 71.33 68.42 43.15 39.88 33.39 6.93 28.76 22.85 2.61
-QSGNN(chunk) - - - 40.18 37.55 13.26 30.25 29.41 8.71
-QSGNN(w/o inter) - - - 40.27 35.98 4.24 28.36 19.19 2.72
-QSGNN(w/o doc) - - - 36.54 32.49 8.01 26.19 26.57 5.02
-QSGNN(w/o pt) 63.66 53.93 36.83 35.64 28.14 5.15 24.62 19.28 2.53
-QSGNN(w/o ft) 77.19 74.69 62.57 51.08 46.42 17.12 43.58 34.23 10.48
-We present the ablation study of QSGNN on Table  10.
-As for the chunk retrieval (QSGNN + chunk, w/o inter or w/o doc), we retrieve the top 10 most relevant chunks for QA and we only report the EM and F1 score since we have no ground truth for chunk retrieval.
-We find that QSGNN(chunk) does not perform as well as QSGNN.
-It can be attributed to two factors: i) QSGNN is not directly trained on chunk labels.
+Quest-GNN 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
+Quest-GNN(chunk) - - - 40.18 37.55 13.26 30.25 29.41 8.71
+Quest-GNN(w/o entity) 60.34 43.31 25.02 27.25 16.56 2.05 19.58 9.65 1.13
+Quest-GNN(w/o chunk) 77.65 74.99 62.60 51.18 45.82 16.55 42.74 35.69 9.08
+Quest-GNN(w/o doc) - - - 36.54 32.49 8.01 26.19 26.57 5.02
+Quest-GNN(w/o intra) 71.33 68.42 43.15 39.88 33.39 6.93 28.76 22.85 2.61
+Quest-GNN(w/o inter) - - - 40.27 35.98 4.24 28.36 19.19 2.72
+Quest-GNN(w/o q_attn) 76.26 71.51 53.74 48.02 40.54 9.65 40.24 30.29 6.56
+Quest-GNN(w/o pt) 63.66 53.93 36.83 35.64 28.14 5.15 24.62 19.28 2.53
+Quest-GNN(w/o ft) 77.19 74.69 62.57 51.08 46.42 17.12 43.58 34.23 10.48
+Table 6. The influence of GNN layer number.
+MuSiQue(Rec@5) MuSiQue(F1)
+Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
+HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24
+Quest-GNN(1 layer) 73.81 69.14 55.13 46.25 41.86 11.46
+Quest-GNN(2 layers) 78.54 76.02 64.65 52.64 47.47 19.03
+Quest-GNN(3 layers) 76.35 74.55 61.66 51.34 45.24 16.45
+Quest-GNN(4 layers) 62.60 58.29 35.37 38.93 29.15 4.61
+Table 7. The influence of negative sampling number.
+MuSiQue(Rec@5) MuSiQue(F1)
+Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
+HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24
+Quest-GNN(Neg_N=1) 68.67 60.27 50.82 43.72 38.25 5.18
+Quest-GNN(Neg_N=10) 77.97 76.16 62.14 51.03 46.68 17.62
+Quest-GNN(Neg_N=30) 78.54 76.02 64.65 52.64 47.47 19.03
+Quest-GNN(Neg_N=50) 77.18 76.69 63.51 51.34 46.99 18.36
+Quest-GNN(Neg_N=100) 77.09 76.16 63.85 52.95 46.37 18.22
+Figure 3. Influence of pre-training scale on various model dimension.
+4.5. Ablation Study
+We present the ablation study of Quest-GNN on Table  5.
+As for the chunk retrieval (Quest-GNN + chunk, w/o inter or w/o doc), we retrieve the top 10 most relevant chunks for QA and we only report the EM and F1 score since we have no ground truth for chunk retrieval.
+Our key findings are:
+  • Quest-GNN(chunk) does not perform as well as Quest-GNN.
+It can be attributed to two factors: i) Quest-GNN is not directly trained on chunk labels.
 ii) Document generally contains more information than chunk.
+  • In terms of the node in Multi-L KG, the model’s performance drops significantly without entities, as they form the basic semantic relationships in Multi-L KG.
+Removing entities causes substantial knowledge loss.
+Removing chunk nodes shows relatively less impact, as document nodes can provide global contextual information to compensate for their absence.
+No document lacks global context information, so it does not perform as well as Quest-GNN(chunk).
+  • In terms of the message passing of Quest-GNN, no intra-level message passing severely deteriorates performance because Quest-GNN loses the ability to directly capture the structure of Multi-L KG but have to be bridged by chunk nodes.
 No inter-level message passing will not update the representations thus have no difference with the embeddings baselines, whose performance decrease a lot when hop number increases.
-No document lacks global context information, so it performs worse than QSGNN(chunk).
-In addition, QSGNN(w/o ft) can achieve competitive performance in 2-hop questions but fail to perform as well as QSGNN, since we pre-train QSGNN only on 1,2-hop questions.
-QSGNN(w/o pt) struggles to perform well because it can not be fitted well on only small number of training pairs.
-A.7 Hyperparameter Discussion
-Here we discussion some of important hyperparameters, including the receptive field (QSGNN layer number), negative sampling number, pre-training data scale and information bottleneck dimension.
-Receptive Field Discussion.
-We discuss the relationships between GNN layers and the multi-hop performance, the results are shown in Table  11.
-We find that 1-layer QSGNN (one intra-level + one inter-level) have limited performance and the gap between 2-layer QSGNN becomes bigger as the hop number increase, it is because 1-layer QSGNN can only aggregate 2-hop information.
-The receptive field limits the potential of QSGNN.
-The 2-layer QSGNN achieves the best performance among 2,3,4 hop questions because it can aggregate more useful information.
+The absence of question alignment has a particularly negative effect on high-hop questions, where the influence of noise will be amplified as the hop number increases.
+  • In terms of the training strategy, Quest-GNN(w/o ft) can achieve competitive performance in 2-hop questions but fail to perform as well as Quest-GNN, since we pre-train Quest-GNN only on 1,2-hop questions.
+Quest-GNN(w/o pt) struggles to perform well because it’s likely to overfit with small number of training pairs.
+4.6. Hyperparameter Discussion
+We discuss some of important hyper-parameters, including the receptive field (Quest-GNN layer number), negative sampling number, pre-training data scale and information bottleneck dimension.
+4.6.1. Receptive Field Discussion
+We discuss the relationships between GNN layers and the multi-hop performance, the results are shown in Table  6.
+We find that 1-layer Quest-GNN (one intra-level + one inter-level) have limited performance and the gap between 2-layer Quest-GNN becomes bigger as the hop number increase, it is because 1-layer Quest-GNN can only aggregate 2-hop information.
+The receptive field limits the potential of Quest-GNN.
+The 2-layer Quest-GNN achieves the best performance among 2,3,4 hop questions because it can aggregate more useful information.
 However, as the layer number increases (4 layers), the performance will suffer a lot because of the over-smoothing problem of GNN  (Rusch et al., 2023; Chen et al., 2020a).
-Table 11: The influence of GNN layer number.
-MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
-Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
-HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
-QSGNN(1 layer) 73.81 69.14 55.13 46.25 41.86 11.46 38.29 29.60 7.83
-QSGNN(2 layers) 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
-QSGNN(3 layers) 76.35 74.55 61.66 51.34 45.24 16.45 42.11 34.50 9.13
-QSGNN(4 layers) 62.60 58.29 35.37 38.93 29.15 4.61 26.68 18.68 2.03
-Negative Sampling Number.
+4.6.2. Negative Sampling Number
 Table 7 shows the impact of negative sampling number on model’s performance.
 According to the results, we find that low sampling number (Neg_N = 1) has limited performance due to inadequate discrimination between target documents and distractions.
-High sampling number (Neg_N=50, 100) shows limited benefits, because excessive easy negatives fail to provide meaningful guidance for QSGNN training.
-Table 12: The influence of negative sampling number.
-MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
-Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
-HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
-QSGNN(Neg_N=1) 68.67 60.27 50.82 43.72 38.25 5.18 34.87 27.44 2.10
-QSGNN(Neg_N=10) 77.97 76.16 62.14 51.03 46.68 17.62 42.28 38.54 10.76
-QSGNN(Neg_N=30) 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
-QSGNN(Neg_N=50) 77.18 76.69 63.51 51.34 46.99 18.36 44.07 38.02 11.65
-QSGNN(Neg_N=100) 77.09 76.16 63.85 52.95 46.37 18.22 43.16 37.52 11.53
-Figure 3: Influence of pre-training scale on various model dimension.
-Pre-training Scale and Information Dimension.
-We pre-train QSGNN on the sythesized QA pairs from 0 to 150k across information dimension from 32 to 512 respectively, the results are shown in Figure  3.
+High sampling number (Neg_N=50, 100) shows limited benefits, because excessive easy negatives fail to provide meaningful guidance for Quest-GNN training.
+4.6.3. Pre-training Scale and Information Dimension
+We pre-train Quest-GNN on the sythesized QA pairs from 0 to 150k across information dimension from 32 to 512 respectively, the results are shown in Figure  3.
 We conduct experiments on the MuSiQue dataset.
 As for the pre-training scale, the results show that insufficient pre-training data leads to sub-optimal performance across all the dimensions.
-As the amount of pre-training data increases, the performance of QSGNN gets better, although the marginal improvement decreases.
-As for the information dimension, the results indicate that increasing the information dimension improves the performance of QSGNN, as larger dimensions enhance model expressiveness through increased model scale.
+As the amount of pre-training data increases, the performance of Quest-GNN gets better, although the marginal improvement decreases.
+As for the information dimension, the results indicate that increasing the information dimension improves the performance of Quest-GNN, as larger dimensions enhance model expressiveness through increased model scale.
 This observation is consistent with the model scaling analysis presented in GFM-RAG  (Luo et al., 2025).
@@
 In this paper, we set the dimension as 128 considering the balance between performance and computational limitation.
-A.8 Good Case Studies
-We present three typical comparative cases between QSGNN and HippoRAG2 to demonstrate QSGNN’s advantages.
-In each case, correct retrieval evidence is marked in green and incorrect evidence in red.
-Notably, QSGNN’s marks are based on query alignment, whereas HippoRAG2’s marks is based on PPR scores  (Gutiérrez et al., 2025).
-Figure 4: Good case 1 for QSGNN.
-Figure 5: Good case 2 for QSGNN.
-Figure 6: Good case 3 for QSGNN.
-Good Case 1.
-Figure  4 illustrates a scenario where QSGNN achieves recall@5 of 1.0 with correct answer generation, while HippoRAG2 fails.
-The 4-hop question requires understanding composite semantics: i) “the only group larger than” corresponds to “the world’s second largest recorded music company”, ii) “the headquarters of” refers to “based in Santa Monica, California” and iii) “the explorer reach the city” denotes “set foot in the area was the party of explorer Gaspar de Portola”.
-QSGNN successfully identifies the critical documents describing Sony Music Entertainment’s ranking, location and the city information, whereas HippoRAG2 erroneously retrieves information about Jive Records’ headquarters.
-This failure stems from HippoRAG2’s selection of seed nodes (“headquarter”, “located in”) without high-level contextual information.
-The case validates Multi-L KG’s semantic modeling capability and query alignment’s effectiveness in filtering misinformation.
-Good Case 2.
-Figure  5 shows QSGNN achieving perfect recall@5 while retrieval of HippoRAG2 is inaccurate.
-For the 4-hop question about Andrew Deveaux’s birthplace, QSGNN integrates evidence across multiple documents and each document describes one key evidence for the question: i) biographical data (“born in South Carolina”), ii) state capital history (“Columbia became capital in 1786”), and iii) geographic relationships (“Forest Acres borders Columbia”).
-The successful retrieval is due to the query-alignment and comprehensive understanding of contextual information.
-However, HippoRAG2 incorrectly retrieves Savannah county information due to over-reliance on “county” and “located on” seed nodes.
-Good Case 3.
-Figure  6 presents a good case of QSGNN where it retrieves correctly and gives the correct answer but HippoRAG2 can’t answer correctly because of the inaccurate retrieval.
-In this case, QSGNN successfully find the snippets for the “where the Yongle Emperor greeted the Karmapa”, “the capitol city of Yaxing Coach’s headquarters location” and the history information about “Yangzhou”, which is the combination results of Multi-L KG and query-specific aggregation.
-HippoRAG2 retrieves redundant Sino-Tibetan relations documents.
-The failure occurs because HippoRAG2 prioritizes high PPR scores for entities “Yongle Emperor” and “Karmapa” without considering their contextual relationships.
-A.9 Bad Case Discussion and Limitations
-We analyze three representative failure cases of QSGNN in document retrieval, each illustrating distinct limitations that inform our future research directions.
-Figure 7: Bad case 1 for QSGNN.
-Figure 8: Bad case 2 for QSGNN.
-Figure 9: Bad case 3 for QSGNN.
-Bad Case 1.
-Figure  7 demonstrates QSGNN’s failure to retrieve documents containing “Freikorps” due to a misspelled query terminology (“free crops” should be “Freikorps”, we changed the “free crops” to “Freikorps” then QSGNN could retrieve correctly).
-This spelling error caused QSGNN to prioritize non-critical entities like “democratic government” and “Germany”.
+5. Conclusion
+In this paper, we propose a novel graph learning method for multi-hop questions retrieval.
+We first design a Multi-information Level Knowledge Graph(Multi-L KG) to model the multi-granular relationships among documents.
+Then we introduce Question-Adaptive Graph Neural Network(Quest-GNN).
+The Quest-GNN aggregates information from multi-levels, together with question alignment, to ensure comprehensive question specific representations while filtering out the noise.
+The pre-training strategy further enhance the ability of Quest-GNN.
+Extensive experiments shows the effectiveness of our method in multi-hop questions.
+Appendix A Appendix
+A.1. Dataset Settings
+Table 8. Dataset statistics
+MuSiQue 2Wiki HotpotQA
+#Pre-train 1-hop QA 91,621 52,122 73,128
+#Pre-train 2-hop QA 58,923 12,097 20,619
+#Fine-tune 2-hop QA 270 782 1000
+#Fine-tune 3-hop QA 482 0 0
+#Fine-tune 4-hop QA 248 218 0
+#Test 2-hop QA 66 172 225
+#Test 3-hop QA 104 0 0
+#Test 4-hop QA 55 53 0
+#Eval 2-hop QA 488 847 1000
+#Eval 3-hop QA 334 0 0
+#Eval 4-hop QA 178 153 0
+The statistics of benchmark datasets are shown in Table  8. For the MuSiQue, 2Wiki and HotpotQA, the extra sample have at least 60%, 60% and 50% supported documents in existing corpora.
+A.2. Baseline Settings
+BM25 is implemented as BM25S (Lù, 2024).
+As for the text embedding based models, we use the NV-Embed-v2-7B and GTE-Qwen2-7B-Instruct from Huggingface  (Wolf et al., 2019).
+For the all the other baselines we use their official implementations.
+For iterative retrievers, we set the max retrieval number as 5.
+For GFM-RAG, we use the official settings to pre-train and fine-tune on our corpora.
+For GNN-RAG we use ReaRev  (Mavromatis and Karypis, 2022) as GNN reasoning.
+All other baseline settings are as their official suggestions.
+A.3. QA Performance on GPT-4o-mini
+Table 9. QA performance on GPT-4o-mini. We highlight the best with bold and the second best with under line.
+MuSiQue 2Wiki HotpotQA
+Method EM F1 EM F1 EM F1
+No Retrieval 10.81 21.39 31.83 36.51 27.44 41.47
+BM25 12.54 23.88 32.25 40.30 34.50 52.62
+Contriever+IRCoT 13.90 25.28 29.44 34.66 30.17 45.55
+GTE-Qwen2-7B 31.15 39.88 50.75 52.37 52.19 66.17
+NV-Embed-v2 28.07 37.15 49.22 51.51 54.85 68.43
+GNN-RAG 23.02 30.12 34.52 48.25 49.12 58.99
+GFM-RAG 28.30 36.81 43.79 53.26 51.86 64.73
+RAPTOR 32.93 41.04 45.43 56.24 55.19 68.16
+GraphRAG 33.31 40.13 50.99 58.08 56.45 69.24
+HippoRAG2 33.12 40.30 51.49 60.18 56.39 70.53
+Quest-GNN(Ours) 34.27 41.78 53.52 61.63 56.04 69.73
+Table 9 presents QA performance using GPT-4o-mini.
+The results show that Quest-GNN achieves competitive EM and F1 scores, demonstrating its compatibility with different LLM backbones.
+Notably, all methods show reduced performance compared to Llama-3-70B-Instruct, primarily due to the intrinsic knowledge gap between GPT-4o-mini and Llama-3.3-70B-Instruct.
+Other observations align with the report of Table  2.
+A.4. Case Study and Discussion
+We present the case study of Quest-GNN here, which includes two good cases and two bad cases.
+Correct evidence is marked in green and incorrect evidence in red.
+Notably, Quest-GNN’s marks are based on question alignment, whereas HippoRAG2’s marks are based on PPR scores  (Gutiérrez et al., 2025).
+Figure 4. Good case 1 for Quest-GNN.
+Figure 5. Good case 2 for Quest-GNN.
+Good Case Analysis.
+Figure  4 and Figure  5 show two good cases where Quest-GNN achieves recall@5 of 1.0 with correct answer generation while retrieval of HippoRAG2 is inaccurate.
+Quest-GNN successfully identifies the critical understanding composite semantics (For example, “the headquarters of” refers to “based in Santa Monica, California” in case 1), and integrates evidence across multiple documents (For example, biographical data:“born in South Carolina”, state capital history:“Columbia became capital in 1786”, and geographic relationships:“Forest Acres borders Columbia” in case 2).
+But HippoRAG2 fails to select the right seed nodes without high-level contextual information.
+These two cases show Multi-L KG’s semantic modeling capability and question alignment’s effectiveness in filtering misinformation and understanding of contextual information comprehensively.
+Figure 6. Bad case 1 for Quest-GNN.
+Figure 7. Bad case 2 for Quest-GNN.
+Bad Case Analysis.
+Figure  6 and Figure  7 reveal two bad cases for Quest-GNN.
+These two cases both stem from poor initial representation.
+In case 1 “Freikorps” is a misspelled query terminology, which should be “free crops” (we changed to the “free crops” then Quest-GNN could retrieve correctly).
+This leads Quest-GNN to prioritize non-critical entities like “democratic government” and “Germany”.
 We also check the pre-training data and fine-tuning data, there is no information about the associations between “free crops” and “Freikorps”, which exacerbates the model’s confusion.
-Bad Case 2.
-Figure  8 reveals QSGNN’s limitation in processing specific term (“San Clemente”, “Fuser”, “Alberto”), which may not be well represented by QSGNN since they are absent from both pre-training and fine-tuning corpora.
-The model deflect to general terms like “province” and “county”, retrieving irrelevant documents.
-Bad Case 3.
-Figure  9 shows a typical bad case for QSGNN.
-In this case the term like “Messi” and “Barcelona” dwarf the key evidence “Diego Maradona, who Messi can bring comparison to”.
-The sequential dependency between finding “Diego Maradona” and subsequent evidence (“June 1982 transfer record”) further complicates retrieval.
-It is difficult for query alignment to identify the correct information because huge amount of documents describing “Messi” and “Barcelona”, which is also related to query, overwhelm the key snippets of “Diego Maradona”.
+In case 2, “San Clemente”, “Fuser” and “Alberto” are all under-represented since they are absent from both pre-training and fine-tuning corpora.
+However, all of them are important for correct retrieval.
 Limitation Discussion and Future Work.
-Based on the analysis of the bad cases above, QSGNN is not a silver bullet for all multi-hop scenarios.
-It exhibits limitations when dealing with specific terminologies especially when they are not well presented in the corpora (see bad case 1, 2), and also struggles to deal the situation where the prevalent mis-related documents overwhelm the key evidence with sequential answer dependencies (see bad case 3).
-We think these limitations can be mitigated by two strategies: i) For the first issue, selecting seed nodes as domain-specific terminologies and sampling constrained subgraphs(depending on the layer of QSGNN, in our case the longest hop is 4), could prioritize relevant documents in retrieval, though this risks information loss.
-ii) For the second issue, integrating Chain-of-Thought(CoT) reasoning  (Chen et al., 2019; Lyu et al., 2023) through query decomposition and iterative retrieval may mitigate the problem.
-However, the first solution is a tricky strategy and even may compromise the performance since subgraph sampling will lead to information loss.
-The second solution introduces CoT into framework, which is beyond the design topic of QSGNN.
-We will leave the combination of these two possible solutions without compromising the performance as the future work of QSGNN.
-A.10 Prompts for OpenIE
-We present the prompts we used for OpenIE workflow.
-Figure  10 shows the prompts used for sentence extraction, Figure  11 shows the prompts for entity extraction, and Figure  12 shows the prompts for triple extraction.
-First, we extract sentences from each document.
-Then we identify core entities from extracted sentences.
-At last, these entities are used to formulate triples from the document.
-Figure 10: Sentence extraction prompt for OpenIE.
-Figure 11: Entity extraction prompt for OpenIE.
-Figure 12: Triple extraction prompt for OpenIE.
-A.11 Prompts for Synthesized Pre-training Data
-The prompts used for generate pre-training data are shown in Figure  13 and Figure  14.
-Figure 13: One hop question generation prompt for OpenIE.
-Figure 14: Two hop question generation prompt for OpenIE.
+Based on the analysis of the bad cases above, Quest-GNN exhibits limitations when dealing with specific terminology or misspelling that can not be well-represented.
+We think this limitation can be mitigated by selecting seed nodes as domain-specific terminologies and sampling constrained subgraphs to prioritize relevant documents in retrieval.
+But how to realize this with limited information loss is still non-trivial.
+We will leave this as our future work.
+A.5. Prompts
+The prompts for OpenIE and question generation are in Figure  8.
+Figure 8. Extraction and question generation prompts.